### v2

# Wenhao Wang

Abstract
This paper explores in-context learning for image copy detection (ICD), i.e., prompting an ICD model to identify replicated images with new tampering patterns without the need for additional training. The prompts (or the contexts) are from a small set of image-replica pairs that reflect the new patterns and are used at inference time. Such in-context ICD has good realistic value, because it requires no fine-tuning and thus facilitates fast reaction against the emergence of unseen patterns. To accommodate the “seen $\rightarrow$ unseen” generalization scenario, we construct the first large-scale pattern dataset named AnyPattern, which has the largest number of tamper patterns ($90$ for training and $10$ for testing) among all the existing ones. We benchmark AnyPattern with popular ICD methods and reveal that existing methods barely generalize to novel patterns. We further propose a simple in-context ICD method named ImageStacker. ImageStacker learns to select the most representative image-replica pairs and employs them as the pattern prompts in a stacking manner (rather than the popular concatenation manner). Experimental results show (1) training with our large-scale dataset substantially benefits pattern generalization ($+26.66\%$ $\mu AP$), (2) the proposed ImageStacker facilitates effective in-context ICD (another round of $+16.75\%$ $\mu AP$), and (3) AnyPattern enables in-context ICD, i.e. without such a large-scale dataset, in-context learning does not emerge even with our ImageStacker. Beyond the ICD task, we also demonstrate how AnyPattern can benefit artists, i.e., the pattern retrieval method trained on AnyPattern can be generalized to identify style mimicry by text-to-image models. The project is publicly available at https://anypattern.github.io.
Figure 1: Top: The comparison between the standard updating process of Image Copy Detection (ICD) and the proposed in-context ICD. Unlike the standard updating approach, our in-context ICD eliminates the need for fine-tuning, making it more efficient. Bottom: AnyPattern is the first large-scale pattern dataset, featuring 90 base and 10 novel patterns. Using 90 base patterns, we generate a training dataset containing 10 million images. Note that each pattern in this paper refers to a class of transformations that are diverse within themselves (see Section A in the Appendix).
1 Introduction
Image Copy Detection (ICD) aims to identify whether a query image is replicated from a database after being tampered with. It serves critical roles in areas such as copyright enforcement, plagiarism prevention, digital forensics, and ensuring content uniqueness on the internet.
Under the realistic scenario, the ICD models suffer from the inevitable emergence of novel tamper patterns. More concretely, the ICD models trained on some already-known patterns may fail when encountering novel patterns. Updating the ICD models for the novel patterns is very expensive and time-consuming. It usually requires collecting a large amount of training samples and then fine-tuning the ICD models, as illustrated in Fig. 1 (a).
As a more efficient solution, this paper explores in-context learning for ICD, as illustrated in Fig. 1 (b). In-context learning is a relatively new machine learning paradigm that learns to solve unseen tasks by providing examples in the prompt. Combining this paradigm with ICD, we endow the ICD models with the ability to recognize novel patterns without fine-tuning. The resulting in-context ICD uses a few examples of image-replica pairs (that prompt the novel patterns) as the context of its input. Though the model parameters remain unchanged, the extracted features are modulated (conditioned) by the context and become competent for recognizing the novel-patterned replication. Consequently, in-context ICD facilitates a fast and efficient reaction against the emergence of unseen patterns.
To set up the “seen $\rightarrow$ unseen” pattern generalization scenario, we construct the first large-scale pattern dataset named AnyPattern. As shown in Fig. 1 (bottom), AnyPattern is featured for its abundant ($100$) tamper patterns, with $90$ for training and $10$ for testing. Concretely, the training set consists of replicas generated from the combination of multiple training patterns (randomly chosen from the $90$ patterns), as well as the original images. We devote approximately one million CPU core hours to generate 10 million training images in total. The testing set consists of queries (replicas) generated from the combination of $10$ novel patterns and galleries (their original images and distractors). Another important characteristic of our AnyPattern dataset is: it provides examples for indicating novel patterns at inference time. Each example is an image-replica pair, within which the replica is generated from the combination of some novel patterns. These examples are very limited (e.g., 10 examples for each pattern combination) and are not to be used for fine-tuning. During inference, the in-context ICD uses these examples to gain knowledge of the novel patterns instantly. A thorough illustration of all the patterns is provided in the Appendix (Section A).
Given the prerequisite dataset, we further propose a simple and straightforward in-context ICD method named ImageStacker. ImageStacker basically follows the standard in-context learning pipeline, i.e., using some examples as the context of the input. The context conditions the feature extraction of the ICD models and is also known as prompt(s). We name our in-context learning method as ImageStacker, because it has a unique prompting manner, i.e., stacking the examples and the input images together. During training, we use the ground truth to prepare the image-replica pairs that have the same patterns as the pseudo query images, yielding the in-context learning. During testing, though we are provided with a set of examples that cover the novel patterns, we still do not know which patterns are exactly the ones for generating the query (the replica). In response, we design a pattern retrieval method to select the image-replica examples that are most likely to share the same patterns with the query images. In other words, we retrieve the most representative image-replica pairs in the example set as the prompts for ImageStacker.
To further demonstrate the significance of introducing the AnyPattern dataset, we present an additional application using AnyPattern with the proposed pattern retrieval method. The text-to-image model can be used to mimic the style of artwork with little cost, and this threatens the livelihoods and creative rights of artists. To help them protect their work, we treat an artist’s ‘style’ as a ‘pattern’ and generalize the trained pattern retrieval method to identify generated images with style mimicry. We provide a detailed discussion of this application of AnyPattern in the Appendix (Section B).
To sum up, this paper makes the following contributions:
  1. 1.
    We introduce in-context ICD, which allows the use of a few examples to prompt an already-trained ICD model to recognize novel-pattern replication, without the need of fine-tuning. To support the scenario of pattern generalization, we construct the first large-scale pattern dataset, i.e. AnyPattern, which provides 90 training patterns and 10 base patterns.
  2. 2.
    We benchmark AnyPattern against (1) popular ICD methods and find that none of these methods could generalize to novel patterns well, and (2) various intuitive prompting approaches, discovering that most of these common visual prompting methods are ineffective for in-context ICD. Therefore, we propose a simple in-context ICD method, i.e. ImageStacker. ImageStacker stacks an image-replica example (prompt) to the query image along the channel dimension, and yields a good in-context learning effect for recognizing novel tamper patterns.
  3. 3.
    To further highlight the generalization and importance of AnyPattern, we present another application of AnyPattern, i.e. the pattern retrieval method trained on AnyPattern can be generalized to identify generated images by text-to-image models that most closely matches the style of a given real artwork.
2 Related Works
In-context Learning for Computer Vision. In-context learning originates from large language models like GPT-3 [1] and Instruction GPT [2]. In the pure computer vision area, in-context learning is a relatively new concept and presents challenges for implementation. The earliest known work, MAE-VQGAN [3], implements in-context learning through image inpainting. Following MAE-VQGAN [3], (Un)SupPR [4] discusses how to find good visual in-context examples; Prompt-SelF [5] explores the factors that affect the performance of visual in-context learning; Painter [6] designs a generalist vision model to automatically generate images according to the example pairs; Prompt Diffusion [7] proposes diffusion-based generative models. Specific computer vision areas also see work such as SegGPT [8] for image segmentation. These methods typically implement in-context learning in a concatenating manner, while we propose a stacking design.
Existing Image Copy Detection Methods. Existing ICD methods [9, 10, 11, 12, 13, 14] can be broadly categorized into contrastive learning-based algorithms and deep metric learning-based algorithms. The use of contrastive learning for training an ICD model is natural due to its reliance on data augmentation. For example, SSCD [9] is based on the InfoNCE [15] loss and introduces a differential entropy regularization to differentiate nearby vectors. CNNCL [10] employs a large memory bank with a contrastive loss to learn from numerous positive and negative pairs. The application of deep metric learning to ICD is intuitive given that ICD is, fundamentally, a retrieval task. EfNet [11] proposes a “drip training” procedure, whereby the number of classes used to train a model is incrementally increased. BoT [12] sets a robust baseline for ICD and introduces descriptor stretching, which normalizes scores at the feature level. However, none of these methods can be directly applied to our in-context ICD scenario as they never consider generalizing their models to novel patterns, let alone using only image-replica pairs.
3 In-context Image Copy Detection
This section first gives a formal definition of in-context Image Copy Detection (ICD) and then introduces the constructed AnyPattern dataset.
3.1 Definition
The in-context ICD requires an already-trained ICD model to recognize novel-pattern replication by using a few image-replica examples as the prompt. As illustrated in Fig. 2, the base tamper patterns underlying the training data have no overlap with the novel patterns, while the prompts are generated from the random combination of novel patterns.
Formally, the objective of in-context ICD is to train a model $g$ with parameters $\tau$ using only the training (base) pattern set. To detect novel-pattern replication, we do not fine-tune the model to update $\tau$, but use some prompts to condition/modify the feature extraction. Given a query image $x_{q}$, its feature is extracted by:
$g_{\tau}\left(\mathcal{F},x_{q}\right),$ (1)
where $\mathcal{F}\subset\mathcal{D}$ are the prompts chosen from the image-replica pool $\mathcal{D}=\left\{\left(A_{i},A^{\prime}_{i}\right)\right\}_{i=1}^{N}$, and $\left(A_{i},A^{\prime}_{i}\right)$ is the $i$th image-replica pair.
In-context ICD requires the extracted feature $g_{\tau}\left(\mathcal{F},x_{q}\right)$ to be discriminative for identifying whether $x_{q}$ is replicated from any gallery image.
Figure 2: The illustration for our in-context Image Copy Detection (ICD) with AnyPattern. In-context ICD necessitates a well-trained ICD model to be prompted to novel patterns with the assistance of a few image-replica pairs and without any fine-tuning process. In realistic scenarios, this setup is highly practical as it provides a feasible solution for a deployed ICD system faced with unseen patterns.
3.2 AnyPattern Dataset
AnyPattern has two characteristics, i.e., 1) having plenty of tamper patterns and 2) providing a small set of image-replica pairs as the prompts of novel patterns.
1) Large size of tamper patterns. AnyPattern set encompasses a total of $100$ patterns: $90$ are designated for training in-context ICD models, while the remaining $10$ are reserved for testing. A comprehensive introduction to these patterns can be found in the Appendix (Section A).
2) A small pool of image-replica pairs. For each combination of novel patterns, we provide $10$ image-replica pairs as the prompts. Totally, there are $1,200$ prompts. We note that during inference, only partial prompts (e.g., $1\sim 10$, depending on the hyper-parameters) are used for each query.
The source images to generate these replicated images and prompts are from the DISC21 dataset [16, 17]. DISC21 has $1$ million unlabeled training images, from which we randomly select $100,000$ images as the un-edited images following [12, 13]. Each training image is transformed $99$ times by randomly selected training patterns. Together with the original training images, we construct a training dataset containing $10$ million images. Owing to the complexity and volume of the patterns, this process is distributed across $200$ CPU nodes in a supercomputing cluster and requires about one million CPU core hours. We adopt the gallery dataset from DISC21 as our gallery. The query set includes $25,000$ queries, among which $5,000$ are generated by applying a randomly selected novel pattern combination to gallery images, and the remaining $20,000$ queries serve as distractors (without true matches in the gallery).
3.3 Comparison against Existing Datasets
Currently, there are three publicly available ICD benchmarks, i.e. CopyDays [18], DISC21 [16, 17], and NDEC [13].
CopyDays [18] was launched in 2009. This dataset contains only 157 query images and 3,000 gallery images, and lacks training data. The types of tampering patterns involved are relatively straightforward, such as alterations in contrast and blurring.
DISC21 [16, 17] was established in 2021 as an extensive benchmark for ICD, notable for its massive scale, including one million training images and one million gallery images, along with complicated tampering patterns. Additionally, it includes numerous distractor queries, which do not correspond to any true matches in the gallery.
NDEC [13] addresses the challenge of hard negatives in ICD, i.e., some images may appear very similar yet are not replications. By incorporating this aspect of hard negatives, NDEC enhances the realism of ICD evaluations.
Beyond these existing datasets, our AnyPattern provides several novel explorations:
(1) AnyPattern has the largest number of tampering patterns. Specifically, our AnyPattern features 90 base patterns and 10 novel patterns. For CopyDays, there are only a few very simple patterns, e.g., contrast changes and blurring. DISC21 features about 20 patterns, including complex ones. NDEC focuses on hard negative problems and inherits the patterns from DISC21.
(2) AnyPattern is the first dataset that carefully regulates the base and novel patterns. All of the previous datasets only define the patterns for generating queries, and none of them restrict the patterns for training. Therefore, researchers directly use the patterns for generating queries as the training patterns, which brings over-optimistic results. In contrast, the test and training patterns in our AnyPattern are well-defined and separated.
(3) AnyPattern is the only dataset that enables in-context ICD. As shown in Table 3, in-context learning does not emerge when training with DISC21 and a small number of patterns. In contrast, with our AnyPattern, training ImageStacker significantly improves performance. This reaffirms the value of our proposed AnyPattern.
Beyond the above discussions, we also introduce the application of AnyPattern to identify style mimicry, highlighting its generalizability and importance. Please refer to Section B in the Appendix.
Figure 3: The proposed ImageStacker includes: (a) prompt selection fetches the most representative image-replica pair from the whole pool for a given query, and (b) prompting design stacks the selected image-replica pair onto a query along the channel dimension, and thus the image-replica pair conditions the feed-forward process. In (c), we show how to unify prompt selection and prompting design into one vision transformer.
4 Method
In this section, we provide a detailed illustration of our proposed ImageStacker (Fig. 3). The deep metric learning baseline used by our ImageStacker is first briefly reviewed in Section 4.1. During testing for queries with novel patterns, though we have a set of examples containing the novel patterns, we still do not know exactly which patterns generate the query (the replica). Hence, in Section 4.2, we propose a prompt selection method, pattern retrieval, to fetch the most representative image-replica pair from the entire image-replica pool. Subsequently, in Section 4.3, using the selected prompt, we introduce a unique prompt design, i.e, stacking. Finally, we try to unify the prompt selection method and the prompting design into one ViT backbone in Section 4.4.
4.1 Baseline
This section briefly overviews the ICD baseline implemented in our ImageStacker. Following [12, 11], we conceptualize ICD as an image retrieval task, primarily adopting deep metric learning methods. Specifically, we treat each original image and all its replicas as a training class and perform deep metric learning on these classes. Pairwise training [19, 20], classification training [21, 22, 23], or their combination can be utilized for this purpose. In our baseline, we select classification training, specifically CosFace [22], due to its demonstrated effectiveness and simplicity.
4.2 Pattern Retrieval
Drawing inspiration from image retrieval techniques, we propose a pattern retrieval method for identifying the image-replica pairs corresponding to a given query (Fig. 3 (a)). The training of pattern retrieval can be seen as a multi-label classification task. When using Vision Transformer (ViT) [24] as the pattern extractor, we design a pattern token $x^{0}_{ptr}$ and concatenate it to the ViT input:
${}\left[x^{L}_{cls},\mathbf{X}^{L},x^{L}_{ptr}\right]=f\left(\left[x^{0}_{cls},\mathbf{X}^{0},x^{0}_{ptr}\right]\right),$ (2)
where $f$ represents the ViT, $\mathbf{X}^{0}$ is the patch tokens, $x^{0}_{cls}$ is the class token, and $L$ is the number of layers in a ViT.
To use $x^{L}_{ptr}$ as the representation of patterns in an image, it is supervised by the binary cross-entropy loss, which is formulated as
$\mathcal{L}_{ptr}=-\frac{1}{M}\sum^{M}_{i=1}\sum^{C}_{c=1}\left[y_{ic}\log(p_{ic})+(1-y_{ic})\log(1-p_{ic})\right],$ (3)
where $M$ is the number of training images, $C$ is the number of training patterns ($C=90$ here), $y_{ic}$ is the label of the $i$-th image for the $c$-th pattern class, and $p_{ic}$ is the predicted probability by the model that the $i$-th image belongs to the $c$-th pattern class. The pattern token interacts with patch tokens during the feed-forward process and can be considered as the feature of a pattern combination. During testing, the classification head is discarded, and retrieval is performed with the feature.
4.3 Stacking
To address the in-context ICD, given a prompt, we introduce a simple yet effective prompting manner, i.e. stacking (Fig. 3 (b)). This unique prompting manner modifies the input structure of a ViT: traditionally, an image is divided into $N$ patches ($\left\{x_{i}\in\mathbb{R}^{\textbf{3}\times P\times P}\mid i=1,2,\ldots,N\right\}$, where $P\times P$ is the patch size) before being passed to the embedding layer. In contrast, ImageStacker stacks an image-replica pair along the channel dimension, tripling the original channel count. As a result, the $N$ patches are represented in a new format:
$\left\{\bar{x_{i}}\in\mathbb{R}^{\textbf{9}\times P\times P}\mid i=1,2,\ldots,N\right\}.$ (4)
To accommodate these $9$-channel image patches, we initialize a new embedding layer (while keeping the hidden size and all other details of the original ViT). For a query/replica, we use the retrieved image-replica pair or ground truth as the prompt; for a gallery/original image, we directly duplicate itself two times as the pseudo-prompt. The stacked image-replica pair alters the feed-forward process, thus allowing for a conditioned input image feature. The advantage of our stacking design lies in introducing an inductive bias to the in-context learning, which emphasizes the contrasts at the same or similar positions between the original image and the copy. Since many tampering patterns occur at the same or similar positions, this inductive bias brings more benefits to the in-context learning process compared to the traditional concatenation method without any inductive bias. In the Appendix (Section C), we further analyze what ViT learns with our ImageStacker design.
4.4 Unifying Pattern Retrieval and Stacking
We try to unify the pattern retrieval and stacking process into one ViT backbone (Fig. 3 (c)). The unification is straightforward for training because the image-replica pair is selected based on ground truths. The final loss is defined as
$\mathcal{L}_{final}=\mathcal{L}_{cls}+\lambda\cdot\mathcal{L}_{ptr},$ (5)
where $\mathcal{L}_{cls}$ is the CosFace loss [22], and $\lambda$ is the balance parameter.
During testing, a key challenge arises: ImageStacker requires an image-replica pair as input; yet before the feed-forward process, we do not have the pattern feature to retrieve an image-replica pair. To overcome this, we introduce the pseudo-image-replica pairs to get the pattern feature for a query, i.e. we duplicate one query itself two times as the pseudo-prompt (Fig. 3 (c-1)). Consequently, we acquire the pattern feature and thus use it to fetch the most representative image-replica pair for each query. Stacking the fetched image-replica pair onto a query, we extract its (image) feature (Fig. 3 (c-2)). Because the gallery does not contain patterns, we duplicate itself two times as its pseudo-prompt and then extract its (image) feature (Fig. 3 (c-3)).
Figure 4: The performance of ICD state-of-the-arts, CLIP [25], and our Basl. (Small) on the DISC21 dataset and novel patterns. Our Basl. (Small) is trained with the same patterns with BoT [12]. Since these algorithms (including our Basl. (Small)) are not designed to handle input of image-replica pairs, we demonstrate their performance decrease by directly testing their trained models on the $10$ novel patterns from AnyPattern.
5 Experiments
5.1 Evaluation Metrics and Training Details
Please refer to the Appendix (Section D and E).
5.2 The Challenge from Novel Patterns
In this section, we present the performance degradation of trained ICD models (SSCD [9], BoT [12], EfNet [11], and CNNCL [10]) and the CLIP [25] model, when encountering novel patterns. All the ICD state-of-the-arts are trained on the DISC21 dataset, and we select the most successful CLIP [25] model, which is trained on the $2$ billion sample English subset of LAION-5B [26] and achieves a zero-shot top-1 accuracy of $80.1\%$ on ImageNet-1k [27]. The corresponding $\mu AP$ and $recall@1$ scores are summarized in Fig. 4, leading us to two main observations. First, all the ICD models, despite being trained with different methods and pattern combinations, experience a significant accuracy decrease (about $60\%$ for $\mu AP$ and $recall@1$) when the evaluation dataset changes from DISC21 to novel patterns. Notably, our baseline is trained on the same pattern combination with BoT [12] (Basl. (Small)) and maintains comparable performance with state-of-the-arts on both the DISC21 dataset and the $10$ novel patterns. Second, while CLIP models display impressive results in zero-shot image classification and image retrieval tasks, they are not ideally suited for ICD tasks (less than $10\%$ $\mu AP$). This can be attributed to CLIP being predominantly trained on natural images. In light of these findings, we argue that in practical scenarios, the continuous emergence of novel patterns poses a significant challenge for deployed ICD models.
Table 1: The performance improvement on the novel patterns from AnyPattern and in-context learning. Please see the performance on base patterns in the Appendix (Section F).
[…]
SmallPattern 14.22 13.72 14.51 13.80 16.18
AnyPattern 39.79 39.81 41.33 40.65 42.84
ImageStacker 53.94 54.11 55.17 53.51 56.65
[…]
SmallPattern 20.24 17.65 21.05 18.02 20.54
AnyPattern 42.83 43.97 47.19 45.68 47.86
ImageStacker 57.31 60.14 59.96 57.10 60.86
Figure 5: The illustration for different visual prompting methods for incorporating image-replica pairs. Designs (2) and (3) operate on the feature level, while designs (4) through (6) function at the image level.
Table 2: The comparison between different visual prompting methods for incorporating image-replica pairs. Our ImageStacker not only achieves the highest performance but also maintains efficient training and inference.
[…]
Basline (Any) $42.84$ $47.86$ 1.28 0.184
PatchIntegrater $37.42$ $42.58$ $4.13$ $0.474$
ClassIntegrater $38.91$ $43.75$ $2.54$ $0.462$
ImageAdder $25.07$ $33.89$ $1.29$ $0.194$
ImageCombiner $48.47$ $53.34$ $5.39$ $0.538$
ImageStacker 56.65 60.86 $1.31$ $0.199$
5.3 The Benefits from AnyPattern and In-context Learning
To improve performance on novel patterns, we show that both directly training models on larger pattern sets and conducting in-context learning are beneficial. As demonstrated in Table 1, we initially train models on the base patterns of our AnyPattern. This expansion significantly enhances performance on novel patterns: for example, resulting in a gain of $26.66\%$ in $\mu AP$ and $27.32\%$ in $recall@1$ for our baseline. However, a noticeable performance gap persists compared to scenarios where training and testing occur on the same pattern sets. This emphasizes the necessity of introducing in-context learning methods to further enhance performance. Our in-context learning method (ImageStacker) not only further improves performance on novel patterns significantly, achieving gains of $+16.75\%$ in $\mu AP$ and $+15.30\%$ in $recall@1$ for our baseline, but also maintains (marginally improves) the performance on base patterns (see Section F in the Appendix).
5.4 Comparing ImageStacker against Common Visual Prompting Methods
Models can integrate image-replica pairs at both feature and image levels. At the feature level, we append an extra self-attention layer to the last block of ViT to enable interaction between the class token of the replica and the patch tokens or class token of the image-replica pairs (see Fig. 5 (2) and (3)). At the image level, besides our proposed ImageStacker, we can directly add the image-replica pairs to the replica (See Fig. 5 (4)) or concatenate the images along the height (width) dimension (see Fig. 5 (5)). We compare these different designs in Table 2, drawing three main observations: (1) Incorporating image-replica pairs at the feature level is ineffective, resulting in about a $4\sim 6\%$ performance drop in $\mu AP$. This is attributable to insufficient interaction between the replica and its image-replica pair. (2) Directly adding image-replica pairs to the replica significantly degrades the performance. This is because imposing a strong priori restriction without specific meaning can be detrimental. (3) Both ImageCombiner and ImageStacker significantly improve performance. Compared to ImageCombiner, ImageStacker achieves a greater performance gain ($+8.18\%$ in $\mu AP$ and $+7.52\%$ in $recall@1$), while requiring only about $1/4$ inference workload and $1/3$ training workload. Also, ImageStacker enhances the baseline by $13.81\%$ in $\mu AP$ and $13.00\%$ in $recall@1$, with nearly the same efficiency.
5.5 AnyPattern Enables In-context ICD
In this section, from a data perspective, we explore factors beyond our proposed ImageStacker that contribute to the emergence of in-context learning. We first adjust the training patterns of ImageStacker from AnyPattern to SmallPattern (the one used in BoT [12]), and subsequently discard the ImageNet-pre-trained models. The performance of the baseline and ImageStacker is presented in Table 3. The Upper Bound is achieved by using the query and its original image as the image-replica (example) during the inference. Our analysis reveals that: data plays a crucial role alongside the model: (1) In-context learning does not emerge when using SmallPattern and ImageNet-pretrained models: comparing against baseline with $16.18\%$ in $\mu AP$, ImageStacker only achieves $15.28\%$ in $\mu AP$. (2) The use of AnyPattern leads to the emergence of in-context learning: even without ImageNet, training ImageStacker on AnyPattern already significantly improves performance ($+13.01\%$ in $\mu AP$ and $+12.35\%$ in $recall@1$). This reaffirms the value of our proposed AnyPattern.
Beyond the above experiments, in the Appendix, we show that (1) the proposed pattern retrieval method is effective (Section G), (2) utilizing multiple image-replica pairs per query further improves the performance (Section I), and (3) our method remains effective even with fewer image-replica pairs in the pool (Section J).
Table 3: The performance of the baseline and our ImageStacker with different (pre-)training data.
[…]
$\checkmark$ $\checkmark$ $42.84$ $47.86$
Baseline $\checkmark$ $\checkmark$ $16.18$ $20.54$
$\checkmark$ $40.52$ $46.16$
$\checkmark$ $\checkmark$ $56.65$ $60.86$
ImageStacker $\checkmark$ $\checkmark$ $15.28$ $20.10$
$\checkmark$ $53.53$ $58.51$
$\checkmark$ $\checkmark$ $99.25$ $99.84$
Upper Bound $\checkmark$ $\checkmark$ $18.11$ $22.83$
$\checkmark$ $99.26$ $99.86$
6 Conclusion
This paper considers a practical scenario, i.e. in-context Image Copy Detection (ICD). Unlike the standard updating of ICD, in-context ICD aims to prompt a trained model to recognize novel-patterned replication using a few example pairs, without requiring re-training. To advance research on in-context ICD, we present AnyPattern, a dataset featuring $100$ tampering patterns. We further propose ImageStacker, a method that directly stacks an example pair onto a query along the channel dimension. The stacking design conditions (modifies) the query feature and thus enables a better matching for the novel patterns. Experimental results highlight the substantial performance improvement gained by AnyPattern and ImageStacker. We hope our work draws research attention to the critical real-world problem in ICD systems, i.e., the fast reaction against novel patterns. Beyond the problem of ICD, we also explore the value of AnyPattern in identifying style mimicry by text-to-image models. Limitations and future work. Although training ImageStacker with AnyPattern significantly improves performance on novel patterns, it still falls short compared to base patterns. Based on AnyPattern, future work may focus on developing more effective and efficient in-context learning methods to reduce overfitting on base patterns and close this performance gap.
Appendix A Demonstration of the AnyPattern dataset
This section shows the $100$ patterns utilized in the creation of the AnyPattern dataset. The original image is presented in Fig. 6; the generated replicas are in the following tables, with the names of the training patterns indicated in black, and the names of the test patterns depicted in blue. The majority of the constructed patterns includes a degree of randomness while a minority of the patterns remain constant. To demonstrate the variability, each pattern is replicated four times for a single image. To our knowledge, the assembled AnyPattern is the most extensive pattern set currently available.
Figure 6: The original image that used to add $100$ different patterns
Pattern Elaboration Demo 1 Demo 2 Demo 3 Demo 4
ResizeCrop
Randomly crop and resize an image to a specified size. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Blend
Blend two images with random transparency. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
GrayScale
Convert an image into grayscale. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
ColorJitter
Randomly change the brightness, contrast, saturation, and hue of an image. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Pattern Elaboration Demo 1 Demo 2 Demo 3 Demo 4
Blur
Randomly apply a blur filter to an image. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Pixelate
Pixelate random portions of an image. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Rotate
Randomly rotate an image within a given range of degrees. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Padding
Pad an image with random colors, height and width. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
AddNoise
Add random noise to an image. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
VertFlip
Flip an image vertically. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
HoriFlip
Flip an image horizontally. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
MemeFormat
Add a random meme onto an image. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Pattern Elaboration Demo 1 Demo 2 Demo 3 Demo 4
OverlayEmoji
Overlay a random Emoji onto an image. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
OverlayText
Overlay some random texts onto an image. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
PerspChange
Randomly transform the perspective of an image. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
OverlayImage
Randomly overlay an image onto another one. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
StackImage
Randomly stack two images along the height or width dimension. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
ChangeChan
Randomly shift, swap, or invert the channel of an image [11]. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
EncQuality
Randomly encode (reduce) the quality of an image. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
AddStripes
Randomly add different colors of stripes onto an image. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Pattern Elaboration Demo 1 Demo 2 Demo 3 Demo 4
Sharpen
Randomly enhance the edge contrast of an image, making it appear more defined and detailed. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Skew
Randomly skew an image by a certain angle. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
ShufPixels
Randomly rearrange (shuffle) the pixels within an image, creating a significant disruption to the original structure. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
AddShapes
Randomly add shapes (such as ellipse, rectangle, triangle, star, and pentagon) onto an image. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Repeat
Randomly repeat an image for several times. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
CutAssemble
Randomly cut an image into several cells and shuffle the cells. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
BGChange
Randomly change the background of an image. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Xraylize
Simulate the effect of an X-ray on an image. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Pattern Elaboration Demo 1 Demo 2 Demo 3 Demo 4
Mirrorize
Reflect an image along its vertical or horizontal axis. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Kaleidoscope
Replicate the effect of a kaleidoscope on an image by creating symmetrical, repeating patterns randomly. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
EdgeDetect
Randomly perform edge detection by different filters. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
GlassEffect
Add glass effect onto an image with random extent. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
OptDistort
Simulate the effect of viewing an image through a medium that randomly distorts the light passing through it. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Weather
Add the weather effect (such as rain, fog, clouds, frost, and snow) onto an image. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
SplitRotate
Randomly split and rotate an image. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Solarize
Invert all pixel values above a random threshold, creating a peculiar high-contrast effect. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Pattern Elaboration Demo 1 Demo 2 Demo 3 Demo 4
Legofy
Randomly make an image look as if it is made out of $1\times 1$ LEGO blocks. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Ink
Randomly make an image look like an ink painting. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Zooming
Randomly simulate the effect of zooming in or out while capturing an image. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
CutPaste
Randomly change two parts in an image. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Cartoonize
Transforms an image into a stylized, simplified version reminiscent of a cartoon. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
FDA
Apply a random Fourier Domain Adaptation [28] to an image. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
OilPaint
Randomly make an image look like an oil painting. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Jigsaw
Randomly make an image look like an Jigsaw. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Pattern Elaboration Demo 1 Demo 2 Demo 3 Demo 4
FlRoCollage
A random combination of the flipped and rotated version of an image. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Elastic
Simulate a jelly-like distortion of an image. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
PatchInter
Dividing two random images into patches and then interweave these patches to create a new composite image. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
FancyPCA
Random alter the intensities of the RGB channels in an image, based on Principal Component Analysis (PCA) of the set of RGB pixel values [29]. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
GridDistort
Apply random non-linear distortions within a grid to an image, producing a variety of warping effects. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
HistMatch
Transform an image so that its histogram matches a random histogram. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
ISONoise
Apply random camera sensor noise to an image. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Multiple
A random value, drawn from a distribution, is multiplied with the pixel values of an image. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Pattern Elaboration Demo 1 Demo 2 Demo 3 Demo 4
Posterize
Randomly reduce the number of bits used to represent the color components of each pixel. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Gamma
Alter the luminance values of an image by applying a power-law function. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Shadowy
Randomly add shadows onto an image. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Gravel
Randomly add gravels onto an image. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
SunFlare
Randomly add sun flares onto an image. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
SpeedUp
Randomly add speeding up effect onto an image. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Season
Randomly change the season of an image to spring, summer, autumn, or winter. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
CorrectExpo
Correct the exposure of an image. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Pattern Elaboration Demo 1 Demo 2 Demo 3 Demo 4
Mosaic
Randomly add mosaic effect on an image. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
PartsCollage
Randomly collage different parts of an image. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
RandCollage
Randomly collage $N\times M$ images. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Spatter
Randomly create a spatter effect on an image. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
SuperPixel
SuperPixel are the random result of perceptual grouping (image over-segmentation) of pixels in an image [30]. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Heatmap
Randomly apply a color map to an image. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
PolarWarp
Randomly apply other patterns in a polar-transformed space. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
DropChannel
Randomly drop a channel of an image. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Pattern Elaboration Demo 1 Demo 2 Demo 3 Demo 4
ColorQuant
Randomly reduce the number of distinct colors used in an image. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Emboss
Replace each pixel of an image with a highlight or shadow to give a three-dimensional impression. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Voronoi
Randomly segment an image based on the ‘closeness’ of pixels to certain features or points. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
ColorSpace
Randomly change an image with RGB space to HSV, HLS, LAB, YCrCb, LUV, XYZ, or YUV space. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
OldSchool
Randomly change an image into old school style. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Ascii
Randomly use colored ASCII texts to represent an image. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
PixelMelt
Randomly simulate the effect of "melting" pixels down the image. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Stylize
Randomly replace the style of an image with another one’s [31]. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Pattern Elaboration Demo 1 Demo 2 Demo 3 Demo 4
Animation
Randomly transfer an image to animation style [32]. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
AdvAttack
Randomly perform adversarial attacks to an image to generate adversarial example [33]. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
CLAHE
Apply Contrast Limited Adaptive Histogram Equalization (CLAHE) [34] to equalize an image. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Screenshot
Emboss an image into a random screen shot. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
DisLocation
Random the location of patches in an image. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
DropArea
Apply DropOut [35] or DropBlock [36] on the image level. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
WaveBlock
Apply WaveBlock [37] on the image level. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Shape
Randomly crop an area with circle, heart, polygon, or star shape from an image. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Pattern Elaboration Demo 1 Demo 2 Demo 3 Demo 4
VanGoghize
Change the style of an image to a painting from Vincent van Gogh randomly. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Moire
Add random Moire patterns onto one image. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Wavelet
Apply Wavelet [38] to an image ans select random frequency as the final result. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Sepia
Change an image to Sepia style. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
ToneCurve
Randomly change the relationship between bright and dark areas of the image by manipulating its tone curve. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Erasing
Erase a random area of an image [39]. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
AutoAug
Apply AutoAugment [40], which is an automated data augmentation technique discovered by the reinforcement learning, to an image. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
AutoSeg
Use the Segment Anything [41] model to preform segmentation on an image automatically. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Pattern Elaboration Demo 1 Demo 2 Demo 3 Demo 4
FGRemove
Remove the foreground of an image. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
ErodeDilate
Randomly add the erosion or dilation effect onto an image. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Fleet
Transform an image into one with a sense of time or the accumulation of years. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
WaterWave
Add random water wave effect onto an image. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Binary
Use different methods to binarize an image. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Dehaze
Remove the haze from an image using dark channel prior [42]. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Pyramid
Stack images to create a pyramid architecture. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Swirl
Create a random swirl effect on an image. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Appendix B AnyPattern helps artists
B.1 Background
DreamBooth [43] presents a novel method for personalizing text-to-image diffusion models. By using a few reference images of a subject, DreamBooth enables the model to generate new, high-quality images of that subject in various contexts specified by textual descriptions. This fine-tuning approach retains the model’s general capabilities while enhancing its ability to produce detailed and contextually appropriate depictions of the specific subject. Due to its potential negative societal impact, with concerns that ‘malicious parties might try to use such images to mislead viewers’, the inventors at Google decided not to release any code or trained models. However, a third party re-implemented it and made it publicly available  [^1] https://github.com/XavierXiao/Dreambooth-Stable-Diffusion.
Figure 7: The demonstration for the style mimicry by a tuned DreamBooth model.
Style mimicry. After the release of DreamBooth, people discover that it can easily be used to mimic the styles of any artist. Specifically, anyone can collect as few as five artworks created by an artist and spend less than one dollar to train a DreamBooth model, which can then generate numerous images in the same style. For instance, Ogbogu Kalu released a tuned DreamBooth model on Hugging Face [^2] https://huggingface.co/ogkalu/Comic-Diffusion, which was tuned using the artworks of six comic artists and has gained significant popularity. In Fig. 7, we contrast the real artworks created by these artists with the images generated by the tuned DreamBooth model, demonstrating that their styles are indeed very similar.
Opinion of artists and others. The phenomenon of style mimicry has sparked a debate about the ethics of fine-tuning AI on the artworks of living artists in the comments on Reddit [^3] https://www.reddit.com/r/StableDiffusion/comments/yaquby/2d_illustration_styles_are_scarce_on_stable/ and other places. Supporters of AIGC consider the generated images ‘incredibly beautiful’ and describe the released model as their ‘favorite custom model by far’. Meanwhile, some legal professionals argue that it is lawful because ‘style is not copyrightable’. However, there is more to it than that. Firstly, artists feel frustrating, uncomfortable and invasive for this [^4] https://waxy.org/2022/11/invasive-diffusion-how-one-unwilling-illustrator-found-herself-turned-into-an-ai-model/. For instance, Hollie Mengert wonders if the model’s creator simply did not think of her as a person. She also has concerns about copyright because some of the training artworks from her are created for copyright holders, such as Disney and Penguin Random House. Secondly, these models may end artists ability to earn a living [44]. Artists invest significant time and effort in cultivating their unique styles, which is a critical aspect of their livelihood. As a model replicates these styles without offering compensation, artists’ opportunities to market their work and connect with potential buyers are significant hindered. Finally, the artistic creation will be a swan song [^5] https://www.kqed.org/arts/13928253/ai-art-artificial-intelligence-student-artists-midjourney. This imitation by AI can be demoralizing for art students who are training to become the next generation of artists. Seeing AI models potentially replace their future roles, even if they successfully carve out their own distinct styles, can be discouraging and impact their career aspirations.
B.2 Our target
Figure 8: The demonstrations for matching the style of given artworks against millions of generated images.
Glaze [44] offers ‘style cloaks’ to artists to help mislead the mimicry of their styles by text-to-image models. However, the authors of Glaze acknowledge a limitation: this preventative approach can only protect newer artworks. More specifically, many works have already been downloaded from art repositories such as ArtStation and DeviantArt, and these artists’ styles can still be mimicked using older artworks collected before Glaze was released.
Therefore, it is important for artists to be aware of any generated images that mimic the styles of their released artworks. By knowing this, they can utilize opt-out and removal options, i.e., requesting that providers of these text-to-image models or the owners of generated images with mimicked styles cease the style mimicry.
Here, we aim to provide artists with such a ‘style retrieval’ tool. It is a direct application of our AnyPattern and pattern retrieval method, demonstrating their generalizability to other datasets or real-world scenarios where patterns significantly differ. Specifically, we treat an artist’s ‘style’ as a ‘pattern’. Therefore, using our pattern retrieval method trained with AnyPattern, we can directly search a database containing millions of generated images to identify the image that most closely matches the style of a given real artwork. A demonstration of such a process is shown in Fig. 8.
B.3 Experiments
Figure 9: Original artworks (left column) with their corresponding top-10 style matches generated by a text-to-image model (right column), showcasing our trained model’s proficiency in capturing color, texture, and thematic elements.
Experimental setup. To test the generalizability of our trained pattern retrieval method in identifying style mimicry by text-to-image models, we first construct a database of millions of generated images. Specifically, we utilize prompts from DiffusionDB [45] and Stable Diffusion V1.5666https://huggingface.co/runwayml/stable-diffusion-v1-5 to generate 1,819,776 images. Then, we collect several publicly available artworks by artists such as Charlie Bowater, Hollie Mengert, Mario Alberti, Pepe Larraz, Andreas Rocha, and James Daly III as query images. It is important to note that our use of these artworks falls under the category of fair use for research purposes [^7] https://en.wikipedia.org/wiki/Fair_use, thereby avoiding any copyright infringement. Employing the style descriptor [^8] https://github.com/WangWenhao0716/AnyPatternStyle we trained on AnyPattern, we extract a 768-dimensional vector from each generated image and real artwork. By computing the cosine similarities, we identify the top-10 generated images that most closely match the style of each real artwork.
Observations. We visualize some matching results in Fig. 9 and conclude that our trained model successfully identifies the style mimicry by text-to-image models: Firstly, the model appears to accurately capture the unique color palettes and lighting of the original artworks. For instance, the vibrant purples and blues in Charlie Bowater’s piece are reflected in the generated images. Similarly, the dusky, sepia tones of Mario Alberti’s cityscapes are well-represented. Secondly, the stylistic elements, like brush strokes and texturing, seem to be well understood by the model. Andreas Rocha’s landscapes with their distinct, somewhat stylized textures are matched with similar generated images. Finally, the thematic elements are reflected in the generated matches. For Hollie Mengert’s character-focused art, the matched generated images also focus on character-centric scenes. Pepe Larraz’s dynamic compositions with a flair for the dramatic are mirrored in the matched images which capture similar energy and movement. James Daly III’s work that features a blend of sci-fi and fantasy elements is matched with images that maintain this blend.
B.4 Limitations and future directions
Although our trained pattern retrieval method successfully generalizes to identify style mimicry, we acknowledge that a gap still exists between the manually-designed patterns in AnyPattern and the art styles created by artists. Therefore, to better assist artists in identifying style mimicry, future work may involve incorporating a broader range of art styles into the training set and developing corresponding quantitative evaluations.
Appendix C Understanding ImageStacker with ViT
This section tries to provide a understanding of ImageStacker with ViT. The ViT can be succinctly described as a series of transformer blocks followed by a Layer Normalization (LN) layer (See Line $76$ to $81$ from vit-pytorch). The following analysis considers a query with three different example pairs: (1) one obtained from pattern retrieval, (2) an incorrect one, serving as the lower bound, and (3) one comprising the query itself coupled with its original image, serving as the upper bound.
Figure 10: The comparison between with and without the last layer normalization.
A sudden change in similarity is observed after the final Layer Normalization layer. Fig. 10 (a) shows the cosine similarities between the features of the query — conditioned by different example pairs — and its original image, following various stages, i.e., the transformer blocks and the LN layer. The displayed cosine similarity represents the average calculation across the novel test set. We observe that: (1) in all three cases, the cosine similarities are consistently high following a series of transformer blocks. This intriguing finding suggests that these features lack discriminative power for the in-context ICD. (2) the final LN layer renders the features discriminative: following this layer, the upper bound ($99.25\%$ $\mu AP$) maintains high similarity ($0.938$), the lower bound ($28.55\%$ $\mu AP$) experiences a sudden decrease in similarity ($0.892$ to $0.307$), and the retrieval one resides between the two bounds. A natural question is what does ViT learn without the final LN layer? To answer this, we remove the last LN layer and conduct the same experiments. Experimental results are shown in Fig. 10 (b): (1) without the last LN layer, the onset of discrimination shifts “forward". Specifically, the cosine similarity begins to change from the $8$th transformer block and exhibits a sharp decrease at the $10$th transformer block; (2) though the performance of our method (Pattern Retrieval) drops by $4.65\%$ without the final LN layer, it is still much higher that the baseline ($42.84\%$), demonstrating the robustness of our ImageStacker with pattern retrieval; (3) even when using the original image in conjunction with the query itself as the example pair (the upper bound), the $\mu AP$ does not remain at $100\%$. This indicates that the emergence of in-context learning is compromised to a certain degree.
Figure 11: The dissection of the final LN layer.
Figure 12: The influence of normalization and affine in the training stage.
Diving into the final Layer Normalization layer. Given the input $x$, the LN is represented by:
$y=\frac{x-\mathrm{E}[x]}{\sqrt{\operatorname{Var}[x]+\epsilon}}*\gamma+\beta,$ (6)
where: $\epsilon$ is a value added to the denominator for numerical stability; $\gamma$ and $\beta$ are two learnable affine transform parameters. It can be dissected into normalization and affine. Fig. 11 shows the change of similarity inside the LN layer. It is the affine transformation that separates different features.
To delve deeper into the roles of normalization and affine transformation during training, we independently remove each: the normalization and the affine transformation. The modified versions of the final LN layer are represented as:
$y_{1}=x*\gamma+\beta,$ (7)
and
$y_{2}=\frac{x-\mathrm{E}[x]}{\sqrt{\operatorname{Var}[x]+\epsilon}}.$ (8)
The repeated experimental results are illustrated in Fig. 12. We discover that: (1) as expected, removing either the normalization or the affine transformation results in a performance drop for our method ($-3.37\%$ and $-1.48\%$ respectively). However, both are still less than the drop observed when the last LN layer is entirely removed ($-4.65\%$); (2) the phenomenon in which discrimination shifts “forward" is attributed to the affine transformation. Specifically, without normalization, there is no “forward" shift, whereas without the affine transformation, a similar phenomenon is observed; (3) even without either the normalization or the affine transformation, the $\mu AP$ upper bound still remains near $100\%$. This indicates that this in-context learning phenomenon is not solely a result of the LN itself; even a basic affine transformation can produce a similar effect.
The final Layer Normalization layer is also crucial for fundamental vision tasks like image classification. This section explores the influence of the final LN layer beyond our in-context ICD. We select the image classification task for demonstration due to its fundamental nature. The experiments are conducted with DeiT [46] on ImageNet [27], with results shown in Fig. 13 and Table 4. We observe that the final LN layer plays a crucial role in (1) stabilizing the training process – without it, the training loss can fluctuate to as high as $35$; and in (2) enhancing the performance – for instance, in the absence of the final LN layer, the top-$1$ accuracy of DeiT-B drops by $2.4\%$ (from $81.8\%$ to $79.4\%$). These findings underscore the significance of the final LN layer, extending beyond the in-context ICD.
Figure 13: The change of training loss of DeiT-B with respect to epochs.
Table 4: Comparison of top-1 accuracy with and without the final LN layer.
Top-1 Accuracy with last LN without last LN
DeiT-Ti $72.2$ $68.4$
DeiT-S $79.8$ $76.1$
DeiT-B $81.8$ $79.4$
Appendix D Evaluation metrics
We employ two evaluation metrics for the in-context ICD, namely $\mu AP$ and $Recall@1$. $\mu AP$ serves as an overall evaluation metric and is equivalent to the area under the Precision-Recall curve, providing a comprehensive measure of both the precision and recall of our model across varying thresholds. On the other hand, the $Recall@1$ metric is query-specific. It checks whether the actual (correct) result appears first in the list of all returned results, offering insight into the effectiveness of our model in accurately retrieving the most relevant result at the top.
Appendix E Training details
We implement our ImageStacker using PyTorch [47] and distribute its training across eight Nvidia A100 GPUs. We use ViT-B/16 [24] as the backbone, which is pre-trained on the ImageNet dataset [27] using DeiT [46] unless otherwise specified. Before training, we resize images to a resolution of $224\times 224$ pixels. We set the balance parameter, $\lambda$, at $1$ and use a batch size of $512$. Each batch adopts the standard PK sampling method, with $128$ classes and $4$ images per class. The total number of training epochs is $25$ with a cosine-decreasing learning rate. The margin $m$ and scale $s$ in CosFace loss [22] are set to $0.35$ and $64$, respectively.
Appendix F The performance on base patterns
Table 5: The performance on the base patterns from AnyPattern and in-context learning. Our in-context learning method (ImageStacker) not only improves performance on novel patterns significantly but also maintains (marginally improves) the performance on base patterns.
[…]
AnyPattern 77.54 76.12 77.56 79.84 79.37
ImageStacker 81.39 79.98 80.11 84.13 83.56
[…]
AnyPattern 81.76 80.37 81.13 83.41 83.05
ImageStacker 84.52 83.96 83.78 86.07 85.85
It is crucial to maintain performance on base patterns while enhancing performance on novel patterns through in-context learning because the edited copies generated by the base patterns may still appear in the future. As illustrated in Table 5, our proposed ImageStacker achieves this objective. For instance, with our baseline, the $\mu AP$ has been increased from $79.37\%$ to $83.56\%$, and the $recall@1$ has improved from $83.05\%$ to $85.85\%$.
Appendix G The proposed pattern retrieval method is effective
Table 6: Different methods to retrieve the image-replica pair of a query.
[…]
Basline (Any) $42.84$ $47.86$
Random $50.42$ $54.83$
Approximate $56.65$ $60.86$
Accurate $56.68$ $60.78$
Ground truth $56.99$ $60.96$
Zero Shot $49.71$ $53.76$
Lower Bound $28.55$ $32.25$
Upper Bound $99.25$ $99.84$
Figure 14: The demonstration of different methods for selecting an image-replica pair. A larger score means a better matching.
Table 6 outlines several methods for obtaining the image-replica pair for a given query: (1) Random - selecting an image-replica pair randomly from the entire image-replica pool. (2) Approximate retrieval - using the ImageStacker model. (3) Accurate retrieval - employing another model. (4) Ground truth - using pattern ground truth of queries (not available in practice). ImageStacker can also be applied in a zero-shot setting by duplicating the query two times as the pseudo-prompt (Zero Shot). The lower bound is obtained by using an incorrect image-replica pair, while the upper bound is achieved by using the query and its original image as the image-replica. Their pattern retrieval accuracy is shown in the Appendix (Section H). The demonstration of these methods is visualized in the Fig. 14 (See more demonstrations in Fig. 17 and Fig. 18). We observe that: (1) our approximate retrieval method achieves performance comparable to that of the accurate method ($-0.03\%$ $\mu AP$), and even comparable to using the pattern ground truth ($-0.34\%$ $\mu AP$). It significantly outperforms the random method ($+6.23\%$ $\mu AP$). (2) The zero-shot setting surpasses the baseline by $6.87\%$ in $\mu AP$ and $5.90\%$ in $recall@1$, demonstrating the generalizability of our method. (3) It is non-trivial that using the query itself and its original image as the image-replica nearly achieves $100\%$ performance (see Table 3: without our AnyPattern, using the query itself and its original image can only improve the $\mu AP$ from $16.18\%$ to $18.11\%$), providing evidence of the emergence of in-context learning again.
Appendix H Our pattern retrieval method achieves high accuracy
Table 7: The accuracy of pattern retrieval methods and corresponding $\mu AP$.
[…]
Random $30.11$ $50.42$
Approximate $62.90$ $56.65$
Accurate $64.60$ $56.68$
Ground truth $100.00$ $56.99$
This section shows the accuracy of our pattern retrieval methods achieve. The accuracy is defined as
$acc\ =\ \frac{1}{N}\sum^{N}_{i=1}acc_{i},$ (9)
$acc_{i}\ =\ \frac{\#\left(P_{q_{i}}\cap P_{s_{i}}\right)}{{}\#\left(P_{q_{i}}\right)},$ (10)
[…]
The pattern retrieval accuracy and the corresponding $\mu AP$ for the four retrieval methods are displayed in Table 7. Our observations are as follows: (1) compared to random selection, our methods of approximate and accurate pattern retrieval show an improvement of approximately $+30\%$ in accuracy, leading to a $+6\%$ improvement in $\mu AP$; (2) while using ground truths can achieve $100\%$ accuracy in pattern retrieval, there is little to no room for further improvement in $\mu AP$. These findings indicate that, although our pattern retrieval is not perfect, it is sufficient to retrieve an example pair to condition the feed-forward process.
Appendix I Utilizing multiple image-replica pairs per query further improves the performance
Figure 15: Leveraging multiple image-replica pairs per query: the change of $\mu AP$ and $recall@1$ in relation to the number of image-replica pairs per query.
We also discover that leveraging multiple image-replica pairs for a given query can further enhance performance. For each query, we retrieve its top-$N$ image-replica pairs and repeat the $9$-channel ImageStacker $N$ times. The final similarity between a query and a gallery is then calculated as the maximum value among $N$ similarities. The performance change in relation to the number of image-replica pairs is depicted in Fig. 15. Compared to using only one image-replica pair, employing $10$ image-replica pairs boosts the $\mu AP$ to $59.59\%$ ($+2.94\%$) and $recall@1$ to $63.16\%$ ($+2.30\%$).
Appendix J Our method remains effective even with fewer image-replica pairs in the pool.
Figure 16: The change of $\mu AP$ and $recall@1$ in relation to the number of image-replica pairs per pattern combination in the pool.
Currently, the AnyPattern pool contains 10 image-replica pairs for each pattern combination. As shown in Fig. 16, we find that for our ImageStacker, one image-replica pair per pattern combination is sufficient to improve performance. This implies that when we utilize only one image-replica pair per query for inference (unlike the scenario described in Section I), 10 image-replica pairs per pattern combination in the pool achieves similar performance to just one pair. This further demonstrates the practicality of our method: when operators of a deployed ICD system identify novel patterns, they need only manually add one example pair to the pool.
Appendix K Author statement
We, the undersigned, hereby affirm that we bear full responsibility for the content provided herein. We guarantee that all data and materials used and presented are compliant with applicable laws and are free from any encroachments on intellectual property rights or any other legal rights of third parties. In the event of any violations, we accept full liability and will address all associated legal consequences.
Appendix L Hosting, licensing, and maintenance plan.
The data is hosted by Hugging Face, Inc. at https://huggingface.co/datasets/WenhaoWang/AnyPattern. The dataset and corresponding codes are released under the MIT License. We commit to maintaining the dataset, which includes regular updates and corrections as necessary.
Figure 17: More visualizations of different methods for selecting an example pair. A larger score means a better matching.
Figure 18: More visualizations of different methods for selecting an example pair. A larger score means a better matching.
Appendix M Data Sheet for AnyPattern
For what purpose was the dataset created? Was there a specific task in mind? Was there a specific gap that needed to be filled? Please provide a description.
This paper explores in-context learning for image copy detection (ICD), i.e., prompting an ICD model to identify replicated images with new tampering patterns without the need for additional training. Unlike the standard updating approach, our in-context ICD eliminates the need for fine-tuning, making it more efficient. To accommodate the “seen → unseen” generalization scenario, we construct the first large-scale pattern dataset named AnyPattern, which has the largest number of tamper patterns (90 for training and 10 for testing) among all the existing ones.
Who created the dataset (e.g., which team, research group) and on behalf of which entity (e.g., company, institution, organization)?
The dataset was created by Wenhao Wang (University of Technology Sydney), Yifan Sun (Baidu Inc.), Zhentao Tan (Peking University and Baidu Inc.), and Yi Yang (Zhejiang University).
Who funded the creation of the dataset? If there is an associated grant, please provide the name of the grantor and the grant name and number.
Funded in part by Faculty of Engineering and Information Technology Scholarship, University of Technology Sydney.
Any other comments?
None.
Composition
What do the instances that comprise the dataset represent (e.g., documents, photos, people, countries)? Are there multiple types of instances (e.g., movies, users, and ratings; people and interactions between them; nodes and edges)? Please provide a description.
Each instance represents an edit copy of an original image or the original image itself.
How many instances are there in total (of each type, if appropriate)?
There are 10,000,000 training images, 1,000,000 reference images, 1,200 image-replica pairs, and 25,000 queries.
Does the dataset contain all possible instances or is it a sample (not necessarily random) of instances from a larger set? If the dataset is a sample, then what is the larger set? Is the sample representative of the larger set (e.g., geographic coverage)? If so, please describe how this representativeness was validated/verified. If it is not representative of the larger set, please describe why not (e.g., to cover a more diverse range of instances, because instances were withheld or unavailable).
The dataset contains all possible instances.
What data does each instance consist of? “Raw” data (e.g., unprocessed text or images)or features? In either case, please provide a description.
Each instance consists of an image.
Is there a label or target associated with each instance? If so, please provide a description.
Yes, each edited copy has a pointer to its original image.
Is any information missing from individual instances? If so, please provide a description, explaining why this information is missing (e.g., because it was unavailable). This does not include intentionally removed information, but might include, e.g., redacted text.
Everything is included. No data is missing.
Are relationships between individual instances made explicit (e.g., users’ movie ratings, social network links)? If so, please describe how these relationships are made explicit.
Not applicable.
Are there recommended data splits (e.g., training, development/validation, testing)? If so, please provide a description of these splits, explaining the rationale behind them.
Yes. We follow the general retrieval task to split training, reference, and queries.
Are there any errors, sources of noise, or redundancies in the dataset? If so, please provide a description.
No. The patterns are all generated by code, and thus there is no error.
Is the dataset self-contained, or does it link to or otherwise rely on external resources (e.g., websites, tweets, other datasets)?
The dataset is entirely self-contained.
Does the dataset contain data that might be considered confidential (e.g., data that is protected by legal privilege or by doctor–patient confidentiality, data that includes the content of individuals’ nonpublic communications)? If so, please provide a description. Unknown to the authors of the datasheet.
No.
Does the dataset contain data that, if viewed directly, might be offensive, insulting, threatening, or might otherwise cause anxiety? If so, please describe why.
No.
Does the dataset identify any subpopulations (e.g., by age, gender)? If so, please describe how these subpopulations are identified and provide a description of their respective distributions within the dataset.
No.
Is it possible to identify individuals (i.e., one or more natural persons), either directly or indirectly (i.e., in combination with other data) from the dataset? If so, please describe how.
No.
Any other comments?
None.
Collection
How was the data associated with each instance acquired? Was the data directly observable (e.g., raw text, movie ratings), reported by subjects (e.g., survey responses), or indirectly inferred/derived from other data (e.g., part-of-speech tags, model-based guesses for age or language)? If the data was reported by subjects or indirectly inferred/derived from other data, was the data validated/verified? If so, please describe how.
The data was auto-generated by the code of each pattern. We release the code at: https://github.com/WangWenhao0716/AnyPattern.
What mechanisms or procedures were used to collect the data (e.g., hardware apparatuses or sensors, manual human curation, software programs, software APIs)? How were these mechanisms or procedures validated?
Not applicable.
If the dataset is a sample from a larger set, what was the sampling strategy (e.g., deterministic, probabilistic with specific sampling probabilities)?
AnyPattern does not sample from a larger set.
Who was involved in the data collection process (e.g., students, crowdworkers, contractors) and how were they compensated (e.g., how much were crowdworkers paid)?
No crowdworkers are needed in the data collection process.
Over what timeframe was the data collected? Does this timeframe match the creation timeframe of the data associated with the instances (e.g., recent crawl of old news articles)? If not, please describe the timeframe in which the data associated with the instances was created.
Not applicable.
Were any ethical review processes conducted (e.g., by an institutional review board)? If so, please provide a description of these review processes, including the outcomes, as well as a link or other access point to any supporting documentation.
There were no ethical review processes conducted.
Did you collect the data from the individuals in question directly, or obtain it via third parties or other sources (e.g., websites)?
The data was auto-generated by the code of each pattern. We release the code at: https://github.com/WangWenhao0716/AnyPattern.
Were the individuals in question notified about the data collection? If so, please describe (or show with screenshots or other information) how notice was provided, and provide a link or other access point to, or otherwise reproduce, the exact language of the notification itself.
Not applicable.
Did the individuals in question consent to the collection and use of their data? If so, please describe (or show with screenshots or other information) how consent was requested and provided, and provide a link or other access point to, or otherwise reproduce, the exact language to which the individuals consented.
Not applicable.
If consent was obtained, were the consenting individuals provided with a mechanism to revoke their consent in the future or for certain uses? If so, please provide a description, as well as a link or other access point to the mechanism (if appropriate).
Not applicable.
Has an analysis of the potential impact of the dataset and its use on data subjects (e.g., a data protection impact analysis) been conducted? If so, please provide a description of this analysis, including the outcomes, as well as a link or other access point to any supporting documentation.
Not applicable.
Any other comments?
None.
Preprocessing
Was any preprocessing/cleaning/labeling of the data done (e.g., discretization or bucketing, tokenization, part-of-speech tagging, SIFT feature extraction, removal of instances, processing of missing values)? If so, please provide a description. If not, you may skip the remaining questions in this section.
Not applicable.
Was the “raw” data saved in addition to the preprocessed/cleaned/labeled data (e.g., to support unanticipated future uses)? If so, please provide a link or other access point to the “raw” data.
Yes, raw data is saved.
Is the software that was used to preprocess/clean/label the data available? If so, please provide a link or other access point.
Not applicable.
Any other comments?
None.
Uses
Has the dataset been used for any tasks already? If so, please provide a description.
No.
Is there a repository that links to any or all papers or systems that use the dataset? If so, please provide a link or other access point.
No.
What (other) tasks could the dataset be used for?
This dataset is specifically designed for in-context image copy detection.
Is there anything about the composition of the dataset or the way it was collected and preprocessed/cleaned/labeled that might impact future uses? For example, is there anything that a dataset consumer might need to know to avoid uses that could result in unfair treatment of individuals or groups (e.g., stereotyping, quality of service issues) or other risks or harms (e.g., legal risks, financial harms)? If so, please provide a description. Is there anything a dataset consumer could do to mitigate these risks or harms?
There is minimal risk for harm: the data were already public. No.
Are there tasks for which the dataset should not be used? If so, please provide a description.
All tasks that utilize this dataset should follow the MIT License.
Any other comments?
 None.
Distribution
Will the dataset be distributed to third parties outside of the entity (e.g., company, institution, organization) on behalf of which the dataset was created? If so, please provide a description.
Yes, the dataset is publicly available on the internet.
How will the dataset will be distributed (e.g., tarball on website, API, GitHub)? Does the dataset have a digital object identifier (DOI)?
The dataset is distributed on the project website: https://anypattern.github.io/. The dataset shares the same DOI as this paper.
When will the dataset be distributed?
The dataset is released in April, 2024.
Will the dataset be distributed under a copyright or other intellectual property (IP) license, and/or under applicable terms of use (ToU)? If so, please describe this license and/or ToU, and provide a link or other access point to, or otherwise reproduce, any relevant licensing terms or ToU, as well as any fees associated with these restrictions.
No.
Have any third parties imposed IP-based or other restrictions on the data associated with the instances? If so, please describe these restrictions, and provide a link or other access point to, or otherwise reproduce, any relevant licensing terms, as well as any fees associated with these restrictions.
No.
Do any export controls or other regulatory restrictions apply to the dataset or to individual instances? If so, please describe these restrictions, and provide a link or other access point to, or otherwise reproduce, any supporting documentation.
No.
Any other comments?
None.
Maintenance
Who will be supporting/hosting/maintaining the dataset?
The authors of this paper will be supporting and maintaining the dataset.
How can the owner/curator/manager of the dataset be contacted (e.g., email address)?
The contact information of the curators of the dataset is listed on the project website: https://anypattern.github.io/.
Is there an erratum? If so, please provide a link or other access point.
There is no erratum for our initial release. Errata will be documented in future releases on the dataset website.
Will the dataset be updated (e.g., to correct labeling errors, add new instances, delete instances)? If so, please describe how often, by whom, and how updates will be communicated to dataset consumers (e.g., mailing list, GitHub)?
Yes, we will monitor cases when users can report harmful images and creators can remove their videos. We may include more patterns in the future.
If the dataset relates to people, are there applicable limits on the retention of the data associated with the instances (e.g., were the individuals in question told that their data would be retained for a fixed period of time and then deleted)? If so, please describe these limits and explain how they will be enforced.
No, this dataset is not related to people.
Will older versions of the dataset continue to be supported/hosted/maintained? If so, please describe how. If not, please describe how its obsolescence will be communicated to dataset consumers.
We will continue to support older versions of the dataset.
If others want to extend/augment/build on/contribute to the dataset, is there a mechanism for them to do so? If so, please provide a description. Will these contributions be validated/verified? If so, please describe how. If not, why not? Is there a process for communicating/distributing these contributions to dataset consumers? If so, please provide a description.
Anyone can extend/augment/build on/contribute to AnyPattern. Potential collaborators can contact the dataset authors.
Any other comments?
None.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: We follow this common practice for writing a paper.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: See the Limitation at the end of our main paper.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory Assumptions and Proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: This is not a theory paper.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental Result Reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: The project (including the proposed dataset AnyPattern and the code for ImageStacker) is publicly available at https://anypattern.github.io under the MIT Licence.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: The project (including the proposed dataset AnyPattern and the code for ImageStacker) is publicly available at https://anypattern.github.io under the MIT Licence.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental Setting/Details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: The experimental settings and details are provided in the main paper as well as the Appendix.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment Statistical Significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [No]
  34. Justification: All the experimental results are significant and stable, and error bars are not reported because it would be too computationally expensive.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments Compute Resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We include the experiments compute resources in each related section.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code Of Ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: We follow the NeurIPS Code of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader Impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: Please see the end of Conclusion and the Data Sheet.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: Our paper poses no such risks.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We cite all the assets used in our paper, and all the corresponding licenses are mentioned and respected.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New Assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: The project (including the proposed dataset AnyPattern and the code for ImageStacker) is publicly available at https://anypattern.github.io under the MIT Licence.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and Research with Human Subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: We do not involve research with human subjects and there is no potential risk.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.

### v3

@@
-Wenhao Wang
+Wenhao Wang    Yifan Sun    Zhentao Tan    Yi Yang
 Abstract
@@
 ImageStacker learns to select the most representative image-replica pairs and employs them as the pattern prompts in a stacking manner (rather than the popular concatenation manner).
-Experimental results show (1) training with our large-scale dataset substantially benefits pattern generalization ($+26.66\%$ $\mu AP$), (2) the proposed ImageStacker facilitates effective in-context ICD (another round of $+16.75\%$ $\mu AP$), and (3) AnyPattern enables in-context ICD, i.e.
-without such a large-scale dataset, in-context learning does not emerge even with our ImageStacker.
+Experimental results show (1) training with our large-scale dataset substantially benefits pattern generalization ($+26.66\%$ $\mu AP$), (2) the proposed ImageStacker facilitates effective in-context ICD (another round of $+16.75\%$ $\mu AP$), and (3) AnyPattern enables in-context ICD, i.e., without such a large-scale dataset, in-context learning does not emerge even with our ImageStacker.
 Beyond the ICD task, we also demonstrate how AnyPattern can benefit artists, i.e., the pattern retrieval method trained on AnyPattern can be generalized to identify style mimicry by text-to-image models.
 The project is publicly available at https://anypattern.github.io.
+Keywords:
+Image Copy Detection AnyPattern In-context Learning Style Mimicry
 Figure 1: Top: The comparison between the standard updating process of Image Copy Detection (ICD) and the proposed in-context ICD.
@@
 Using 90 base patterns, we generate a training dataset containing 10 million images.
-Note that each pattern in this paper refers to a class of transformations that are diverse within themselves (see Section A in the Appendix).
+Note that each pattern in this paper refers to a class of transformations that are diverse within themselves (see Appendix (Section A)).
 1 Introduction
@@
 To further demonstrate the significance of introducing the AnyPattern dataset, we present an additional application using AnyPattern with the proposed pattern retrieval method.
-The text-to-image model can be used to mimic the style of artwork with little cost, and this threatens the livelihoods and creative rights of artists.
+The text-to-image diffusion model can be used to mimic the style of artwork with little cost, and this threatens the livelihoods and creative rights of artists.
 To help them protect their work, we treat an artist’s ‘style’ as a ‘pattern’ and generalize the trained pattern retrieval method to identify generated images with style mimicry.
-We provide a detailed discussion of this application of AnyPattern in the Appendix (Section B).
 To sum up, this paper makes the following contributions:
@@
   3. 3.
-    To further highlight the generalization and importance of AnyPattern, we present another application of AnyPattern, i.e.
-the pattern retrieval method trained on AnyPattern can be generalized to identify generated images by text-to-image models that most closely matches the style of a given real artwork.
+    To further highlight the generalization and importance of AnyPattern dataset, we present its another application, i.e.
+the pattern retrieval method trained on AnyPattern can be generalized to identify generated images by text-to-image diffusion models that most closely matches the style of a given real artwork.
 2 Related Works
 In-context Learning for Computer Vision.
-In-context learning originates from large language models like GPT-3 [1] and Instruction GPT [2].
+In-context learning originates from large language models like GPT-3 (Brown et al., 2020) and Instruction GPT (Ouyang et al., 2022).
 In the pure computer vision area, in-context learning is a relatively new concept and presents challenges for implementation.
-The earliest known work, MAE-VQGAN [3], implements in-context learning through image inpainting.
-Following MAE-VQGAN [3], (Un)SupPR [4] discusses how to find good visual in-context examples; Prompt-SelF [5] explores the factors that affect the performance of visual in-context learning; Painter [6] designs a generalist vision model to automatically generate images according to the example pairs; Prompt Diffusion [7] proposes diffusion-based generative models.
-Specific computer vision areas also see work such as SegGPT [8] for image segmentation.
+The earliest known work, MAE-VQGAN (Bar et al., 2022), implements in-context learning through image inpainting.
+Following MAE-VQGAN (Bar et al., 2022), (Un)SupPR (Zhang et al., 2023) discusses how to find good visual in-context examples; Prompt-SelF (Sun et al., 2023) explores the factors that affect the performance of visual in-context learning; Painter (Wang et al., 2023b) designs a generalist vision model to automatically generate images according to the example pairs; Prompt Diffusion (Wang et al., 2023d) proposes diffusion-based generative models.
+Specific computer vision areas also see work such as SegGPT (Wang et al., 2023c) for image segmentation.
 These methods typically implement in-context learning in a concatenating manner, while we propose a stacking design.
 Existing Image Copy Detection Methods.
-Existing ICD methods [9, 10, 11, 12, 13, 14] can be broadly categorized into contrastive learning-based algorithms and deep metric learning-based algorithms.
+Existing ICD methods (Pizzi et al., 2022; Yokoo, 2021; Papadakis and Addicam, 2021; Wang et al., 2021, 2023a; Fernandez et al., 2023) can be broadly categorized into contrastive learning-based algorithms and deep metric learning-based algorithms.
 The use of contrastive learning for training an ICD model is natural due to its reliance on data augmentation.
-For example, SSCD [9] is based on the InfoNCE [15] loss and introduces a differential entropy regularization to differentiate nearby vectors.
-CNNCL [10] employs a large memory bank with a contrastive loss to learn from numerous positive and negative pairs.
+For example, SSCD (Pizzi et al., 2022) is based on the InfoNCE (Oord et al., 2018) loss and introduces a differential entropy regularization to differentiate nearby vectors.
+CNNCL (Yokoo, 2021) employs a large memory bank with a contrastive loss to learn from numerous positive and negative pairs.
 The application of deep metric learning to ICD is intuitive given that ICD is, fundamentally, a retrieval task.
-EfNet [11] proposes a “drip training” procedure, whereby the number of classes used to train a model is incrementally increased.
-BoT [12] sets a robust baseline for ICD and introduces descriptor stretching, which normalizes scores at the feature level.
+EfNet (Papadakis and Addicam, 2021) proposes a “drip training” procedure, whereby the number of classes used to train a model is incrementally increased.
+BoT (Wang et al., 2021) sets a robust baseline for ICD and introduces descriptor stretching, which normalizes scores at the feature level.
 However, none of these methods can be directly applied to our in-context ICD scenario as they never consider generalizing their models to novel patterns, let alone using only image-replica pairs.
@@
 We note that during inference, only partial prompts (e.g., $1\sim 10$, depending on the hyper-parameters) are used for each query.
-The source images to generate these replicated images and prompts are from the DISC21 dataset [16, 17].
-DISC21 has $1$ million unlabeled training images, from which we randomly select $100,000$ images as the un-edited images following [12, 13].
+The source images to generate these replicated images and prompts are from the DISC21 dataset (Douze et al., 2021; Papakipos et al., 2022).
+DISC21 has $1$ million unlabeled training images, from which we randomly select $100,000$ images as the un-edited images following (Wang et al., 2021, 2023a).
 Each training image is transformed $99$ times by randomly selected training patterns.
@@
 3.3 Comparison against Existing Datasets
-Currently, there are three publicly available ICD benchmarks, i.e. CopyDays [18], DISC21 [16, 17], and NDEC [13].
-CopyDays [18] was launched in 2009.
+Currently, there are three publicly available ICD benchmarks, i.e. CopyDays (Douze et al., 2009), DISC21 (Douze et al., 2021), and NDEC (Wang et al., 2023a).
+CopyDays (Douze et al., 2009) was launched in 2009.
 This dataset contains only 157 query images and 3,000 gallery images, and lacks training data.
 The types of tampering patterns involved are relatively straightforward, such as alterations in contrast and blurring.
-DISC21 [16, 17] was established in 2021 as an extensive benchmark for ICD, notable for its massive scale, including one million training images and one million gallery images, along with complicated tampering patterns.
+DISC21 (Douze et al., 2021; Papakipos et al., 2022) was established in 2021 as an extensive benchmark for ICD, notable for its massive scale, including one million training images and one million gallery images, along with complicated tampering patterns.
 Additionally, it includes numerous distractor queries, which do not correspond to any true matches in the gallery.
-NDEC [13] addresses the challenge of hard negatives in ICD, i.e., some images may appear very similar yet are not replications.
+NDEC (Wang et al., 2023a) addresses the challenge of hard negatives in ICD, i.e., some images may appear very similar yet are not replications.
 By incorporating this aspect of hard negatives, NDEC enhances the realism of ICD evaluations.
@@
 This reaffirms the value of our proposed AnyPattern.
-Beyond the above discussions, we also introduce the application of AnyPattern to identify style mimicry, highlighting its generalizability and importance. Please refer to Section B in the Appendix.
 Figure 3: The proposed ImageStacker includes: (a) prompt selection fetches the most representative image-replica pair from the whole pool for a given query, and (b) prompting design stacks the selected image-replica pair onto a query along the channel dimension, and thus the image-replica pair conditions the feed-forward process.
@@
 This section briefly overviews the ICD baseline implemented in our ImageStacker.
-Following [12, 11], we conceptualize ICD as an image retrieval task, primarily adopting deep metric learning methods.
+Following (Wang et al., 2021; Papadakis and Addicam, 2021), we conceptualize ICD as an image retrieval task, primarily adopting deep metric learning methods.
 Specifically, we treat each original image and all its replicas as a training class and perform deep metric learning on these classes.
-Pairwise training [19, 20], classification training [21, 22, 23], or their combination can be utilized for this purpose.
-In our baseline, we select classification training, specifically CosFace [22], due to its demonstrated effectiveness and simplicity.
+Pairwise training (Sohn, 2016; Hermans et al., 2017), classification training (Liu et al., 2016; Wang et al., 2018; Sun et al., 2020), or their combination can be utilized for this purpose.
+In our baseline, we select classification training, specifically CosFace (Wang et al., 2018), due to its demonstrated effectiveness and simplicity.
 4.2 Pattern Retrieval
@@
 The training of pattern retrieval can be seen as a multi-label classification task.
-When using Vision Transformer (ViT) [24] as the pattern extractor, we design a pattern token $x^{0}_{ptr}$ and concatenate it to the ViT input:
+When using Vision Transformer (ViT) (Dosovitskiy et al., 2020) as the pattern extractor, we design a pattern token $x^{0}_{ptr}$ and concatenate it to the ViT input:
 ${}\left[x^{L}_{cls},\mathbf{X}^{L},x^{L}_{ptr}\right]=f\left(\left[x^{0}_{cls},\mathbf{X}^{0},x^{0}_{ptr}\right]\right),$ (2)
@@
 During testing, the classification head is discarded, and retrieval is performed with the feature.
+Figure 4: The demonstration for the style mimicry by a tuned DreamBooth model.
 4.3 Stacking
@@
 Since many tampering patterns occur at the same or similar positions, this inductive bias brings more benefits to the in-context learning process compared to the traditional concatenation method without any inductive bias.
-In the Appendix (Section C), we further analyze what ViT learns with our ImageStacker design.
 4.4 Unifying Pattern Retrieval and Stacking
@@
 $\mathcal{L}_{final}=\mathcal{L}_{cls}+\lambda\cdot\mathcal{L}_{ptr},$ (5)
-where $\mathcal{L}_{cls}$ is the CosFace loss [22], and $\lambda$ is the balance parameter.
+where $\mathcal{L}_{cls}$ is the CosFace loss (Wang et al., 2018), and $\lambda$ is the balance parameter.
+Figure 5: The demonstrations for matching the style of given artworks against millions of generated images.
 During testing, a key challenge arises: ImageStacker requires an image-replica pair as input; yet before the feed-forward process, we do not have the pattern feature to retrieve an image-replica pair.
@@
 3 (c-3)).
-Figure 4: The performance of ICD state-of-the-arts, CLIP [25], and our Basl.
+5 AnyPattern helps artists
+5.1 Background
+DreamBooth (Ruiz et al., 2023) presents a novel method for personalizing text-to-image diffusion models.
+By using a few reference images of a subject, DreamBooth enables the model to generate new, high-quality images of that subject in various contexts specified by textual descriptions.
+This fine-tuning approach retains the model’s general capabilities while enhancing its ability to produce detailed and contextually appropriate depictions of the specific subject.
+Due to its potential negative societal impact, with concerns that ‘malicious parties might try to use such images to mislead viewers’, the inventors at Google decided not to release any code or trained models.
+However, a third party re-implemented it and made it publicly available (Xiao, 2022).
+Style mimicry.
+After the release of DreamBooth, people discover that it can easily be used to mimic the styles of any artist.
+Specifically, anyone can collect as few as five artworks created by an artist and spend less than one dollar to train a DreamBooth model, which can then generate numerous images in the same style.
+For instance, Ogbogu Kalu released a tuned DreamBooth model on Hugging Face (ogkalu, 2024), which was tuned using the artworks of six comic artists and has gained significant popularity.
+In Fig.
+4, we contrast the real artworks created by these artists with the images generated by the tuned DreamBooth model, demonstrating that their styles are indeed very similar.
+Opinion of artists and others.
+The phenomenon of style mimicry has sparked a debate about the ethics of fine-tuning AI on the artworks of living artists in the comments on Reddit (user, 2022) and other places.
+Supporters of AIGC consider the generated images ‘incredibly beautiful’ and describe the released model as their ‘favorite custom model by far’.
+Meanwhile, some legal professionals argue that it is lawful because ‘style is not copyrightable’.
+However, there is more to it than that.
+Firstly, artists feel frustrating, uncomfortable and invasive for this (Baio, 2022).
+For instance, Hollie Mengert wonders if the model’s creator simply did not think of her as a person.
+She also has concerns about copyright because some of the training artworks from her are created for copyright holders, such as Disney and Penguin Random House.
+Secondly, these models may end artists ability to earn a living (Shan et al., 2023).
+Artists invest significant time and effort in cultivating their unique styles, which is a critical aspect of their livelihood.
+As a model replicates these styles without offering compensation, artists’ opportunities to market their work and connect with potential buyers are significant hindered.
+Finally, the artistic creation will be a swan song (Nguyen, 2023).
+This imitation by AI can be demoralizing for art students who are training to become the next generation of artists.
+Seeing AI models potentially replace their future roles can be discouraging and impact their career aspirations.
+5.2 Our target
+Glaze (Shan et al., 2023) offers ‘style cloaks’ to artists to help mislead the mimicry of their styles by text-to-image models.
+However, the authors of Glaze acknowledge a limitation: this preventative approach can only protect newer artworks.
+More specifically, many works have already been downloaded from art repositories such as ArtStation and DeviantArt, and these artists’ styles can still be mimicked using older artworks collected before Glaze was released.
+Therefore, it is important for artists to be aware of any generated images that mimic the styles of their released artworks.
+By knowing this, they can utilize opt-out and removal options, i.e., requesting that providers of these text-to-image models or the owners of generated images with mimicked styles cease the style mimicry.
+Here, we aim to provide artists with such a ‘style retrieval’ tool.
+It is a direct application of our AnyPattern and pattern retrieval method, showing their generalizability to other datasets or real-world scenarios where patterns significantly differ.
+Specifically, we treat an artist’s ‘style’ as a ‘pattern’.
+Therefore, using our pattern retrieval method trained with AnyPattern, we can directly search a database containing millions of generated images to identify the image that most closely matches the style of a given real artwork.
+A demonstration of such a process is shown in Fig.
+5.
+5.3 Implementation
+Figure 6: Original artworks (left column) with their corresponding top-10 style matches generated by a text-to-image model (right column), showcasing our trained model’s proficiency in capturing color, texture, and thematic elements.
+Experimental setup.
+To test the generalizability of our trained pattern retrieval method in identifying style mimicry by text-to-image models, we first construct a database of millions of generated images.
+Specifically, we utilize prompts from DiffusionDB (Wang et al., 2023e) and Stable Diffusion V1.5 (RunwayML, 2022) to generate $1,819,776$ images.
+Then, we collect several publicly available artworks by artists such as Charlie Bowater, Hollie Mengert, Mario Alberti, Pepe Larraz, Andreas Rocha, and James Daly III as query images.
+It is important to note that our use of these artworks falls under the category of fair use for research purposes (contributors, 2024), thereby avoiding any copyright infringement.
+Employing the style descriptor [^1] https://github.com/WangWenhao0716/AnyPatternStyle we trained on AnyPattern, we extract a 768-dimensional vector from each generated image and real artwork.
+By computing the cosine similarities, we identify the top-10 generated images that most closely match the style of each real artwork.
+Observations.
+We visualize some matching results in Fig.
+6 and conclude that our trained model successfully identifies the style mimicry by text-to-image models: Firstly, the model appears to accurately capture the unique color palettes and lighting of the original artworks.
+For instance, the vibrant purples and blues in Charlie Bowater’s piece are reflected in the generated images.
+Similarly, the dusky, sepia tones of Mario Alberti’s cityscapes are well-represented.
+Secondly, the stylistic elements, like brush strokes and texturing, seem to be well understood by the model.
+Andreas Rocha’s landscapes with their distinct, somewhat stylized textures are matched with similar generated images.
+Finally, the thematic elements are reflected in the generated matches.
+For Hollie Mengert’s character-focused art, the matched generated images also focus on character-centric scenes.
+Pepe Larraz’s dynamic compositions with a flair for the dramatic are mirrored in the matched images which capture similar energy and movement.
+James Daly III’s work that features a blend of sci-fi and fantasy elements is matched with images that maintain this blend.
+5.4 Limitations and future directions
+Although our trained pattern retrieval method successfully generalizes to identify style mimicry, we acknowledge that a gap still exists between the manually-designed patterns in AnyPattern and the art styles created by artists.
+Therefore, to better assist artists in identifying style mimicry, future work may involve incorporating a broader range of art styles into the training set and developing corresponding quantitative evaluations.
+Figure 7: The performance of ICD state-of-the-arts, CLIP (Radford et al., 2021), and our Basl.
 (Small) on the DISC21 dataset and novel patterns.
 Our Basl.
-(Small) is trained with the same patterns with BoT [12].
+(Small) is trained with the same patterns with BoT (Wang et al., 2021).
 Since these algorithms (including our Basl.
@@
 6.1 Evaluation Metrics and Training Details
-Please refer to the Appendix (Section D and E).
+Evaluation metrics.
+We employ two evaluation metrics for the in-context ICD, namely $\mu AP$ and $recall@1$/$R@1$.
+$\mu AP$ serves as an overall evaluation metric and is equivalent to the area under the Precision-Recall curve, providing a comprehensive measure of both the precision and recall of our model across varying thresholds.
+On the other hand, the $recall@1$ metric is query-specific.
+It checks whether the actual (correct) result appears first in the list of all returned results, offering insight into the effectiveness of our model in accurately retrieving the most relevant result at the top.
+Table 1: The performance improvement on the novel patterns from AnyPattern and in-context learning.
+[…]
+SSCD
+BoT
+EfNet
+CNNCL
+Baseline
+SmallPattern $14.22$ $13.72$ $14.51$ $13.80$ $16.18$
+AnyPattern $39.79$ $39.81$ $41.33$ $40.65$ $42.84$
+ImageStacker $53.94$ $54.11$ $55.17$ $53.51$ $\mathbf{56.65}$
+[…]
+SSCD
+BoT
+EfNet
+CNNCL
+Baseline
+SmallPattern $20.24$ $17.65$ $21.05$ $18.02$ $20.54$
+AnyPattern $42.83$ $43.97$ $47.19$ $45.68$ $47.86$
+ImageStacker $57.31$ $60.14$ $59.96$ $57.10$ $\mathbf{60.86}$
+Training details.
+We implement our ImageStacker using PyTorch (Paszke et al., 2019) and distribute its training across eight Nvidia A100 GPUs.
+We use ViT-B/16 (Dosovitskiy et al., 2020) as the backbone, which is pre-trained on the ImageNet dataset (Deng et al., 2009) using DeiT (Touvron et al., 2021) unless otherwise specified.
+Before training, we resize images to a resolution of $224\times 224$ pixels.
+We set the balance parameter, $\lambda$, at $1$ and use a batch size of $512$.
+Each batch adopts the standard PK sampling method, with $128$ classes and $4$ images per class.
+The total number of training epochs is $25$ with a cosine-decreasing learning rate.
+The margin $m$ and scale $s$ in CosFace loss Wang et al.
+(2018) are set to $0.35$ and $64$, respectively.
 6.2 The Challenge from Novel Patterns
-In this section, we present the performance degradation of trained ICD models (SSCD [9], BoT [12], EfNet [11], and CNNCL [10]) and the CLIP [25] model, when encountering novel patterns.
-All the ICD state-of-the-arts are trained on the DISC21 dataset, and we select the most successful CLIP [25] model, which is trained on the $2$ billion sample English subset of LAION-5B [26] and achieves a zero-shot top-1 accuracy of $80.1\%$ on ImageNet-1k [27].
+In this section, we present the performance degradation of trained ICD models (SSCD (Pizzi et al., 2022), BoT (Wang et al., 2021), EfNet (Papadakis and Addicam, 2021), and CNNCL (Yokoo, 2021)) and the CLIP (Radford et al., 2021) model, when encountering novel patterns.
+All the ICD state-of-the-arts are trained on the DISC21 dataset, and we select the most successful CLIP (Radford et al., 2021) model, which is trained on the $2$ billion sample English subset of LAION-5B (Schuhmann et al., 2022) and achieves a zero-shot top-1 accuracy of $80.1\%$ on ImageNet-1k (Deng et al., 2009).
 The corresponding $\mu AP$ and $recall@1$ scores are summarized in Fig.
-4, leading us to two main observations.
+7, leading us to two main observations.
 First, all the ICD models, despite being trained with different methods and pattern combinations, experience a significant accuracy decrease (about $60\%$ for $\mu AP$ and $recall@1$) when the evaluation dataset changes from DISC21 to novel patterns.
-Notably, our baseline is trained on the same pattern combination with BoT [12] (Basl.
+Notably, our baseline is trained on the same pattern combination with BoT (Wang et al., 2021) (Basl.
 (Small)) and maintains comparable performance with state-of-the-arts on both the DISC21 dataset and the $10$ novel patterns.
@@
 In light of these findings, we argue that in practical scenarios, the continuous emergence of novel patterns poses a significant challenge for deployed ICD models.
-Table 1: The performance improvement on the novel patterns from AnyPattern and in-context learning. Please see the performance on base patterns in the Appendix (Section F).
+Table 2: The performance on the base patterns from AnyPattern and in-context learning.
+Our in-context learning method (ImageStacker) not only improves performance on novel patterns significantly but also maintains (marginally improves) the performance on base patterns.
 […]
-SmallPattern 14.22 13.72 14.51 13.80 16.18
-AnyPattern 39.79 39.81 41.33 40.65 42.84
-ImageStacker 53.94 54.11 55.17 53.51 56.65
+SSCD
+BoT
+EfNet
+CNNCL
+Baseline
+AnyPattern $77.54$ $76.12$ $77.56$ $79.84$ $79.37$
+ImageStacker $81.39$ $79.98$ $80.11$ $84.13$ $83.56$
 […]
-SmallPattern 20.24 17.65 21.05 18.02 20.54
-AnyPattern 42.83 43.97 47.19 45.68 47.86
-ImageStacker 57.31 60.14 59.96 57.10 60.86
+SSCD
+BoT
+EfNet
+CNNCL
+Baseline
+AnyPattern $81.76$ $80.37$ $81.13$ $83.41$ $83.05$
+ImageStacker $84.52$ $83.96$ $83.78$ $86.07$ $85.85$
 Figure 8: The illustration for different visual prompting methods for incorporating image-replica pairs.
 Designs (2) and (3) operate on the feature level, while designs (4) through (6) function at the image level.
+6.3 The Benefits from AnyPattern and In-context Learning
+To improve performance on novel patterns, we show that both directly training models on larger pattern sets and conducting in-context learning are beneficial.
+As demonstrated in Table 1, we initially train models on the base patterns of our AnyPattern.
+This expansion significantly enhances performance on novel patterns: for example, resulting in a gain of $26.66\%$ in $\mu AP$ and $27.32\%$ in $recall@1$ for our baseline.
+However, a noticeable performance gap persists compared to scenarios where training and testing occur on the same pattern sets.
+This emphasizes the necessity of introducing in-context learning methods to further enhance performance.
+Our in-context learning method (ImageStacker) further improves performance on novel patterns significantly, achieving gains of $+16.75\%$ in $\mu AP$ and $+15.30\%$ in $recall@1$ for our baseline.
+Furthermore, it is also crucial to maintain performance on base patterns while enhancing performance on novel patterns through in-context learning because the edited copies generated by the base patterns may still appear in the future.
+As illustrated in Table 2, our proposed ImageStacker achieves this objective.
+For instance, with our baseline, the $\mu AP$ has been increased from $79.37\%$ to $83.56\%$, and the $recall@1$ has improved from $83.05\%$ to $85.85\%$.
+Table 3: The performance of the baseline and our ImageStacker with different (pre-)training data. Without large-scale AnyPattern, in-context learning does not emerge.
+  Method
+AnyPattern
+SmallPattern
+ImageNet
+$\mu AP$ $R@1$
+$\checkmark$ $\checkmark$ $42.84$ $47.86$
+Baseline $\checkmark$ $\checkmark$ $16.18$ $20.54$
+$\checkmark$ $40.52$ $46.16$
+$\checkmark$ $\checkmark$ $\mathbf{56.65}$ $\mathbf{60.86}$
+Image- $\checkmark$ $\checkmark$ $15.28$ $20.10$
+Stacker $\checkmark$ $53.53$ $58.51$
+$\checkmark$ $\checkmark$ $99.25$ $99.84$
+Upper $\checkmark$ $\checkmark$ $18.11$ $22.83$
+Bound $\checkmark$ $99.26$ $99.86$
+6.4 AnyPattern Enables In-context ICD
+In this section, from a data perspective, we explore factors beyond our proposed ImageStacker that contribute to the emergence of in-context learning.
+We first adjust the training patterns of ImageStacker from AnyPattern to SmallPattern (the one used in BoT (Wang et al., 2021)), and subsequently discard the ImageNet-pre-trained models.
+The performance of the baseline and ImageStacker is presented in Table 3.
+The Upper Bound is achieved by using the query and its original image as the image-replica (example) during the inference.
+Our analysis reveals that: data plays a crucial role alongside the model: (1) In-context learning does not emerge when using SmallPattern and ImageNet-pretrained models: comparing against baseline with $16.18\%$ in $\mu AP$, ImageStacker only achieves $15.28\%$ in $\mu AP$.
+(2) The use of AnyPattern leads to the emergence of in-context learning: even without ImageNet, training ImageStacker on AnyPattern already significantly improves performance ($+13.01\%$ in $\mu AP$ and $+12.35\%$ in $recall@1$).
+This reaffirms the value of our proposed AnyPattern dataset.
+6.5 In-context Learning Surpasses Fine-tuning
+Beyond the efficiency advantage, in-context learning also offers two performance advantages, as show in Table 4: (1) Fine-tuning fails when the amount of training data is limited.
+For instance, when using the $1,200$ image-replica pairs ($2,400$ images) in AnyPattern for fine-tuning instead of in-context learning, the $\mu AP$ on novel patterns is $41.15\%$, which is $-15.50\%$ compared to our in-context solution; and (2) Fine-tuning on large-scale data generated by novel patterns can lead to catastrophic forgetting of the base patterns.
+For instance, when using $200,000$ images generated by novel patterns for fine-tuning, compared to the in-context solution, although there is a performance ($+10.87\%$ $\mu AP$) superiority on novel patterns, the performance on base patterns drops significantly ($-63.5\%$ $\mu AP$).
+Table 4: The performance comparison between the in-context solution against the fine-tuning one.
+$2,400$, $20,000$, and $200,000$ represent the number of images used for in-context learning and fine-tuning, respectively.
+In-context
+Fine-tuning
+$\mu AP$ $2,400$
+$2,400$
+$20,000$
+$200,000$
+Novel patterns $56.65$ $41.15$ $60.01$ $67.52$
+Base patterns $83.56$ $73.44$ $73.25$ $20.06$
+In-context
+Fine-tuning
+$R@1$ $2,400$
+$2,400$
+$20,000$
+$200,000$
+Novel patterns $60.86$ $46.70$ $65.73$ $77.44$
+Base patterns $85.85$ $78.04$ $77.77$ $27.99$
+6.6 Ablation Studies
 Table 5: The comparison between different visual prompting methods for incorporating image-replica pairs.
 Our ImageStacker not only achieves the highest performance but also maintains efficient training and inference.
-[…]
+‘Infer’ and ‘Train’ are in ‘$10^{-3}s/img$’ and ‘$s/iter$’, respectively.
+  Method
+$\mu AP$
+$R@1$
+Infer
+Train
 Basline (Any) $42.84$ $47.86$ 1.28 0.184
@@
 ImageStacker 56.65 60.86 $1.31$ $0.199$
-5.3 The Benefits from AnyPattern and In-context Learning
-To improve performance on novel patterns, we show that both directly training models on larger pattern sets and conducting in-context learning are beneficial.
-As demonstrated in Table 1, we initially train models on the base patterns of our AnyPattern.
-This expansion significantly enhances performance on novel patterns: for example, resulting in a gain of $26.66\%$ in $\mu AP$ and $27.32\%$ in $recall@1$ for our baseline.
-However, a noticeable performance gap persists compared to scenarios where training and testing occur on the same pattern sets.
-This emphasizes the necessity of introducing in-context learning methods to further enhance performance.
-Our in-context learning method (ImageStacker) not only further improves performance on novel patterns significantly, achieving gains of $+16.75\%$ in $\mu AP$ and $+15.30\%$ in $recall@1$ for our baseline, but also maintains (marginally improves) the performance on base patterns (see Section F in the Appendix).
-5.4 Comparing ImageStacker against Common Visual Prompting Methods
+Figure 9: The demonstration of different methods for selecting an image-replica pair. A larger score means a better matching.
+ImageStacker outperforms common visual prompting methods.
 Models can integrate image-replica pairs at both feature and image levels.
@@
 Also, ImageStacker enhances the baseline by $13.81\%$ in $\mu AP$ and $13.00\%$ in $recall@1$, with nearly the same efficiency.
-5.5 AnyPattern Enables In-context ICD
-In this section, from a data perspective, we explore factors beyond our proposed ImageStacker that contribute to the emergence of in-context learning.
-We first adjust the training patterns of ImageStacker from AnyPattern to SmallPattern (the one used in BoT [12]), and subsequently discard the ImageNet-pre-trained models.
-The performance of the baseline and ImageStacker is presented in Table 3.
-The Upper Bound is achieved by using the query and its original image as the image-replica (example) during the inference.
-Our analysis reveals that: data plays a crucial role alongside the model: (1) In-context learning does not emerge when using SmallPattern and ImageNet-pretrained models: comparing against baseline with $16.18\%$ in $\mu AP$, ImageStacker only achieves $15.28\%$ in $\mu AP$.
-(2) The use of AnyPattern leads to the emergence of in-context learning: even without ImageNet, training ImageStacker on AnyPattern already significantly improves performance ($+13.01\%$ in $\mu AP$ and $+12.35\%$ in $recall@1$).
-This reaffirms the value of our proposed AnyPattern.
-Beyond the above experiments, in the Appendix, we show that (1) the proposed pattern retrieval method is effective (Section G), (2) utilizing multiple image-replica pairs per query further improves the performance (Section I), and (3) our method remains effective even with fewer image-replica pairs in the pool (Section J).
-Table 3: The performance of the baseline and our ImageStacker with different (pre-)training data.
+Table 6: Different methods to retrieve the image-replica pair of a query.
+  Method $\mu AP$ $R@1$ Pattern Acc.
+Basline (Any) $42.84$ $47.86$
+Random $50.42$ $54.83$ $30.11$
+Approximate $56.65$ $60.86$ $62.90$
+Accurate $56.68$ $60.78$ $64.60$
+Ground truth $56.99$ $60.96$ $100.00$
+Zero Shot $49.71$ $53.76$
+Lower Bound $28.55$ $32.25$
+Upper Bound $99.25$ $99.84$
+The proposed pattern retrieval method is effective.
+Table 6 outlines several methods for obtaining the image-replica pair for a given query: (1) Random - selecting an image-replica pair randomly from the entire image-replica pool.
+(2) Approximate retrieval - using the ImageStacker model.
+(3) Accurate retrieval - employing another model.
+(4) Ground truth - using pattern ground truth of queries (not available in practice).
+ImageStacker can also be applied in a zero-shot setting by duplicating the query two times as the pseudo-prompt (Zero Shot).
+The lower bound is obtained by using an incorrect image-replica pair, while the upper bound is achieved by using the query and its original image as the image-replica.
+Their pattern retrieval accuracy is shown in the next section.
+The demonstration of these methods is visualized in the Fig.
+9.
+We observe that: (1) our approximate retrieval method achieves performance comparable to that of the accurate method ($-0.03\%$ $\mu AP$), and even comparable to using the pattern ground truth ($-0.34\%$ $\mu AP$).
+It significantly outperforms the random method ($+6.23\%$ $\mu AP$).
+(2) The zero-shot setting surpasses the baseline by $6.87\%$ in $\mu AP$ and $5.90\%$ in $recall@1$, demonstrating the generalizability of our method.
+(3) It is non-trivial that using the query itself and its original image as the image-replica nearly achieves $100\%$ performance (see Table 3: without our AnyPattern, using the query itself and its original image can only improve the $\mu AP$ from $16.18\%$ to $18.11\%$), providing evidence of the emergence of in-context learning again.
+Figure 10: Leveraging multiple image-replica pairs per query: the change of $\mu AP$ and $recall@1$ in relation to the number of image-replica pairs per query.
+Figure 11: The change of $\mu AP$ and $recall@1$ in relation to the number of image-replica pairs per pattern combination in the pool.
+Our pattern retrieval method achieves high accuracy. This section shows the accuracy of our pattern retrieval methods achieve. The accuracy is defined as
+$acc\ =\ \frac{1}{N}\sum_{i=1}^{N}acc_{i}=\frac{1}{N}\sum_{i=1}^{N}\left(\frac{\#\left(P_{q_{i}}\cap P_{s_{i}}\right)}{{}\#\left(P_{q_{i}}\right)}\right),$ (6)
 […]
-$\checkmark$ $\checkmark$ $42.84$ $47.86$
-Baseline $\checkmark$ $\checkmark$ $16.18$ $20.54$
-$\checkmark$ $40.52$ $46.16$
-$\checkmark$ $\checkmark$ $56.65$ $60.86$
-ImageStacker $\checkmark$ $\checkmark$ $15.28$ $20.10$
-$\checkmark$ $53.53$ $58.51$
-$\checkmark$ $\checkmark$ $99.25$ $99.84$
-Upper Bound $\checkmark$ $\checkmark$ $18.11$ $22.83$
-$\checkmark$ $99.26$ $99.86$
+The pattern retrieval accuracy and the corresponding performance for the four retrieval methods are displayed in Table 6.
+Our observations are as follows: (1) compared to random selection, our method of approximate pattern retrieval shows an improvement of $+32.79\%$ in $acc$, leading to a $+6.23\%$ improvement in $\mu AP$; (2) while using ground truths can achieve $100\%$ accuracy in pattern retrieval, there is little to no room for further improvement in $\mu AP$.
+These findings indicate that, although our pattern retrieval is not perfect, it is relatively sufficient to retrieve an example pair to condition the feed-forward process.
+Figure 12: The failure cases of our method include: (a) the example pair is incorrect, as in case (1), where the query contains the Pyramid pattern, while the example contains Mosaic and AutoSeg patterns; and (b) the presence of numerous visual distractors, as in case (4), where some references depict scenarios that are inherently similar to the original image.
+The original image is highlighted in green, while incorrect matches are indicated in red.
+Utilizing multiple image-replica pairs per query further improves the performance.
+We also discover that leveraging multiple image-replica pairs for a given query can further enhance performance.
+For each query, we retrieve its top-$N$ image-replica pairs and repeat the $9$-channel ImageStacker $N$ times.
+The final similarity between a query and a gallery is then calculated as the maximum value among $N$ similarities.
+The performance change in relation to the number of image-replica pairs is depicted in Fig.
+10.
+Compared to using only one image-replica pair, employing $10$ image-replica pairs boosts the $\mu AP$ to $59.59\%$ ($+2.94\%$) and $recall@1$ to $63.16\%$ ($+2.30\%$).
+Our method remains effective even with fewer image-replica pairs in the pool.
+Currently, the AnyPattern pool contains 10 image-replica pairs for each pattern combination.
+As shown in Fig.
+11, we find that for our ImageStacker, one image-replica pair per pattern combination is sufficient to improve performance.
+This implies that when we utilize only one image-replica pair per query for inference, 10 image-replica pairs per pattern combination in the pool achieves similar performance to just one pair.
+This further demonstrates the practicality of our method: when operators of a deployed ICD system identify novel patterns, they need only manually add one example pair to the pool.
+6.7 Failure Cases
+As shown in Fig.
+12, we conclude two failure cases: (1) Our method may fail when it retrieves incorrect example pairs.
+This is reasonable because in-context learning fundamentally relies on the example pairs to “learn”; and (2) Our method may fail when there are many visual distractors.
+The hard negative problem is indeed a longstanding issue for ICD.
+To further enhance performance, future work could focus on addressing these two types of failure cases.
 7 Conclusion
@@
 We hope our work draws research attention to the critical real-world problem in ICD systems, i.e., the fast reaction against novel patterns.
-Beyond the problem of ICD, we also explore the value of AnyPattern in identifying style mimicry by text-to-image models.
+Beyond the problem of ICD, we also explore the value of AnyPattern in identifying style mimicry by text-to-image diffusion models.
 Limitations and future work.
@@
 Based on AnyPattern, future work may focus on developing more effective and efficient in-context learning methods to reduce overfitting on base patterns and close this performance gap.
+Appendix
 Appendix A Demonstration of the AnyPattern dataset
@@
 The original image is presented in Fig.
-6; the generated replicas are in the following tables, with the names of the training patterns indicated in black, and the names of the test patterns depicted in blue.
+13; the generated replicas are in the following tables, with the names of the training patterns indicated in black, and the names of the test patterns depicted in blue.
 The majority of the constructed patterns includes a degree of randomness while a minority of the patterns remain constant.
@@
 To our knowledge, the assembled AnyPattern is the most extensive pattern set currently available.
-Figure 6: The original image that used to add $100$ different patterns
+The online version of this demonstration is available at https://huggingface.co/datasets/WenhaoWang/AnyPattern/viewer.
            Pattern Elaboration Demo 1 Demo 2 Demo 3 Demo 4
@@
 Randomly change the brightness, contrast, saturation, and hue of an image. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
+Figure 13: The original image that used to add $100$ different patterns
            Pattern Elaboration Demo 1 Demo 2 Demo 3 Demo 4
@@
 ChangeChan
-Randomly shift, swap, or invert the channel of an image [11]. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
+Randomly shift, swap, or invert the channel of an image (Papadakis and Addicam, 2021). [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
 EncQuality
@@
 Sharpen
-Randomly enhance the edge contrast of an image, making it appear more defined and detailed. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
+Randomly enhance the edge contrast of an image. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
 Skew
@@
 ShufPixels
-Randomly rearrange (shuffle) the pixels within an image, creating a significant disruption to the original structure. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
+Randomly rearrange (shuffle) the pixels within an image. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
 AddShapes
@@
 Kaleidoscope
-Replicate the effect of a kaleidoscope on an image by creating symmetrical, repeating patterns randomly. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
+Replicate the effect of a kaleidoscope on an image. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
 EdgeDetect
@@
 OptDistort
-Simulate the effect of viewing an image through a medium that randomly distorts the light passing through it. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
+Simulate the effect of viewing an image through a medium that randomly distorts the light. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
 Weather
@@
 Zooming
-Randomly simulate the effect of zooming in or out while capturing an image. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
+Randomly simulate the effect of zooming in or out. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
 CutPaste
@@
 FDA
-Apply a random Fourier Domain Adaptation [28] to an image. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
+Apply a random Fourier Domain Adaptation (Yang and Soatto, 2020) to an image. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
 OilPaint
@@
 PatchInter
-Dividing two random images into patches and then interweave these patches to create a new composite image. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
+Dividing two random images into patches and then interweave them. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
 FancyPCA
-Random alter the intensities of the RGB channels in an image, based on Principal Component Analysis (PCA) of the set of RGB pixel values [29].
-[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
+Random use PCA (Krizhevsky et al., 2012) to alter the intensities of the RGB channels in an image. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
 GridDistort
-Apply random non-linear distortions within a grid to an image, producing a variety of warping effects. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
+Apply random non-linear distortions within a grid to an image. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
 HistMatch
@@
 SuperPixel
-SuperPixel are the random result of perceptual grouping (image over-segmentation) of pixels in an image [30]. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
+Random perceptual group pixels in an image (Neubert and Protzel, 2014). [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
 Heatmap
@@
 ColorSpace
-Randomly change an image with RGB space to HSV, HLS, LAB, YCrCb, LUV, XYZ, or YUV space. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
+Randomly change an image with RGB space to HSV, HLS, LAB, YCrCb, LUV, XYZ, or space. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
 OldSchool
@@
 PixelMelt
-Randomly simulate the effect of "melting" pixels down the image. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
+Randomly simulate the effect of ”melting” pixels down the image. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
 Stylize
-Randomly replace the style of an image with another one’s [31]. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
+Randomly replace the style of an image with another one’s (Zhang and Dana, 2017). [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
            Pattern Elaboration Demo 1 Demo 2 Demo 3 Demo 4
 Animation
-Randomly transfer an image to animation style [32]. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
+Randomly transfer an image to animation style (Chen et al., 2020). [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
 AdvAttack
-Randomly perform adversarial attacks to an image to generate adversarial example [33]. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
+Randomly perform adversarial attacks to an image (Kim, 2020). [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
 CLAHE
-Apply Contrast Limited Adaptive Histogram Equalization (CLAHE) [34] to equalize an image. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
+Apply Contrast Limited Adaptive Histogram Equalization (Reza, 2004) to equalize an image. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
 Screenshot
@@
 DropArea
-Apply DropOut [35] or DropBlock [36] on the image level. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
+Apply DropOut (Srivastava et al., 2014) or DropBlock (Ghiasi et al., 2018). [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
 WaveBlock
-Apply WaveBlock [37] on the image level. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
+Apply WaveBlock (Wang et al., 2022) on the image level. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
 Shape
@@
 Wavelet
-Apply Wavelet [38] to an image ans select random frequency as the final result. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
+Apply Wavelet (Torrence and Compo, 1998) with random frequency to an image. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
 Sepia
@@
 ToneCurve
-Randomly change the relationship between bright and dark areas of the image by manipulating its tone curve. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
+Randomly change the bright and dark areas of the image by manipulating its tone curve. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
 Erasing
-Erase a random area of an image [39]. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
+Erase a random area of an image (Zhong et al., 2020). [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
 AutoAug
-Apply AutoAugment [40], which is an automated data augmentation technique discovered by the reinforcement learning, to an image.
-[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
+Apply AutoAugment (Cubuk et al., 2018) to an image. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
 AutoSeg
-Use the Segment Anything [41] model to preform segmentation on an image automatically. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
+Use the Segment Anything (Kirillov et al., 2023) model to preform segmentation. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
            Pattern Elaboration Demo 1 Demo 2 Demo 3 Demo 4
@@
 Dehaze
-Remove the haze from an image using dark channel prior [42]. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
+Remove the haze from an image using dark channel prior (He et al., 2010). [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
 Pyramid
@@
 Create a random swirl effect on an image. [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
-Appendix B AnyPattern helps artists
-B.1 Background
-DreamBooth [43] presents a novel method for personalizing text-to-image diffusion models.
-By using a few reference images of a subject, DreamBooth enables the model to generate new, high-quality images of that subject in various contexts specified by textual descriptions.
-This fine-tuning approach retains the model’s general capabilities while enhancing its ability to produce detailed and contextually appropriate depictions of the specific subject.
-Due to its potential negative societal impact, with concerns that ‘malicious parties might try to use such images to mislead viewers’, the inventors at Google decided not to release any code or trained models.
-However, a third party re-implemented it and made it publicly available  [^1] https://github.com/XavierXiao/Dreambooth-Stable-Diffusion.
-Figure 7: The demonstration for the style mimicry by a tuned DreamBooth model.
-Style mimicry.
-After the release of DreamBooth, people discover that it can easily be used to mimic the styles of any artist.
-Specifically, anyone can collect as few as five artworks created by an artist and spend less than one dollar to train a DreamBooth model, which can then generate numerous images in the same style.
-For instance, Ogbogu Kalu released a tuned DreamBooth model on Hugging Face [^2] https://huggingface.co/ogkalu/Comic-Diffusion, which was tuned using the artworks of six comic artists and has gained significant popularity.
-In Fig.
-7, we contrast the real artworks created by these artists with the images generated by the tuned DreamBooth model, demonstrating that their styles are indeed very similar.
-Opinion of artists and others.
-The phenomenon of style mimicry has sparked a debate about the ethics of fine-tuning AI on the artworks of living artists in the comments on Reddit [^3] https://www.reddit.com/r/StableDiffusion/comments/yaquby/2d_illustration_styles_are_scarce_on_stable/ and other places.
-Supporters of AIGC consider the generated images ‘incredibly beautiful’ and describe the released model as their ‘favorite custom model by far’.
-Meanwhile, some legal professionals argue that it is lawful because ‘style is not copyrightable’.
-However, there is more to it than that.
-Firstly, artists feel frustrating, uncomfortable and invasive for this [^4] https://waxy.org/2022/11/invasive-diffusion-how-one-unwilling-illustrator-found-herself-turned-into-an-ai-model/.
-For instance, Hollie Mengert wonders if the model’s creator simply did not think of her as a person.
-She also has concerns about copyright because some of the training artworks from her are created for copyright holders, such as Disney and Penguin Random House.
-Secondly, these models may end artists ability to earn a living [44].
-Artists invest significant time and effort in cultivating their unique styles, which is a critical aspect of their livelihood.
-As a model replicates these styles without offering compensation, artists’ opportunities to market their work and connect with potential buyers are significant hindered.
-Finally, the artistic creation will be a swan song [^5] https://www.kqed.org/arts/13928253/ai-art-artificial-intelligence-student-artists-midjourney.
-This imitation by AI can be demoralizing for art students who are training to become the next generation of artists.
-Seeing AI models potentially replace their future roles, even if they successfully carve out their own distinct styles, can be discouraging and impact their career aspirations.
-B.2 Our target
-Figure 8: The demonstrations for matching the style of given artworks against millions of generated images.
-Glaze [44] offers ‘style cloaks’ to artists to help mislead the mimicry of their styles by text-to-image models.
-However, the authors of Glaze acknowledge a limitation: this preventative approach can only protect newer artworks.
-More specifically, many works have already been downloaded from art repositories such as ArtStation and DeviantArt, and these artists’ styles can still be mimicked using older artworks collected before Glaze was released.
-Therefore, it is important for artists to be aware of any generated images that mimic the styles of their released artworks.
-By knowing this, they can utilize opt-out and removal options, i.e., requesting that providers of these text-to-image models or the owners of generated images with mimicked styles cease the style mimicry.
-Here, we aim to provide artists with such a ‘style retrieval’ tool.
-It is a direct application of our AnyPattern and pattern retrieval method, demonstrating their generalizability to other datasets or real-world scenarios where patterns significantly differ.
-Specifically, we treat an artist’s ‘style’ as a ‘pattern’.
-Therefore, using our pattern retrieval method trained with AnyPattern, we can directly search a database containing millions of generated images to identify the image that most closely matches the style of a given real artwork.
-A demonstration of such a process is shown in Fig.
-8.
-B.3 Experiments
-Figure 9: Original artworks (left column) with their corresponding top-10 style matches generated by a text-to-image model (right column), showcasing our trained model’s proficiency in capturing color, texture, and thematic elements.
-Experimental setup.
-To test the generalizability of our trained pattern retrieval method in identifying style mimicry by text-to-image models, we first construct a database of millions of generated images.
-Specifically, we utilize prompts from DiffusionDB [45] and Stable Diffusion V1.5666https://huggingface.co/runwayml/stable-diffusion-v1-5 to generate 1,819,776 images.
-Then, we collect several publicly available artworks by artists such as Charlie Bowater, Hollie Mengert, Mario Alberti, Pepe Larraz, Andreas Rocha, and James Daly III as query images.
-It is important to note that our use of these artworks falls under the category of fair use for research purposes [^7] https://en.wikipedia.org/wiki/Fair_use, thereby avoiding any copyright infringement.
-Employing the style descriptor [^8] https://github.com/WangWenhao0716/AnyPatternStyle we trained on AnyPattern, we extract a 768-dimensional vector from each generated image and real artwork.
-By computing the cosine similarities, we identify the top-10 generated images that most closely match the style of each real artwork.
-Observations.
-We visualize some matching results in Fig.
-9 and conclude that our trained model successfully identifies the style mimicry by text-to-image models: Firstly, the model appears to accurately capture the unique color palettes and lighting of the original artworks.
-For instance, the vibrant purples and blues in Charlie Bowater’s piece are reflected in the generated images.
-Similarly, the dusky, sepia tones of Mario Alberti’s cityscapes are well-represented.
-Secondly, the stylistic elements, like brush strokes and texturing, seem to be well understood by the model.
-Andreas Rocha’s landscapes with their distinct, somewhat stylized textures are matched with similar generated images.
-Finally, the thematic elements are reflected in the generated matches.
-For Hollie Mengert’s character-focused art, the matched generated images also focus on character-centric scenes.
-Pepe Larraz’s dynamic compositions with a flair for the dramatic are mirrored in the matched images which capture similar energy and movement.
-James Daly III’s work that features a blend of sci-fi and fantasy elements is matched with images that maintain this blend.
-B.4 Limitations and future directions
-Although our trained pattern retrieval method successfully generalizes to identify style mimicry, we acknowledge that a gap still exists between the manually-designed patterns in AnyPattern and the art styles created by artists.
-Therefore, to better assist artists in identifying style mimicry, future work may involve incorporating a broader range of art styles into the training set and developing corresponding quantitative evaluations.
-Appendix C Understanding ImageStacker with ViT
-This section tries to provide a understanding of ImageStacker with ViT.
-The ViT can be succinctly described as a series of transformer blocks followed by a Layer Normalization (LN) layer (See Line $76$ to $81$ from vit-pytorch).
-The following analysis considers a query with three different example pairs: (1) one obtained from pattern retrieval, (2) an incorrect one, serving as the lower bound, and (3) one comprising the query itself coupled with its original image, serving as the upper bound.
-Figure 10: The comparison between with and without the last layer normalization.
-A sudden change in similarity is observed after the final Layer Normalization layer.
-Fig.
-10 (a) shows the cosine similarities between the features of the query — conditioned by different example pairs — and its original image, following various stages, i.e., the transformer blocks and the LN layer.
-The displayed cosine similarity represents the average calculation across the novel test set.
-We observe that: (1) in all three cases, the cosine similarities are consistently high following a series of transformer blocks.
-This intriguing finding suggests that these features lack discriminative power for the in-context ICD.
-(2) the final LN layer renders the features discriminative: following this layer, the upper bound ($99.25\%$ $\mu AP$) maintains high similarity ($0.938$), the lower bound ($28.55\%$ $\mu AP$) experiences a sudden decrease in similarity ($0.892$ to $0.307$), and the retrieval one resides between the two bounds.
-A natural question is what does ViT learn without the final LN layer?
-To answer this, we remove the last LN layer and conduct the same experiments.
-Experimental results are shown in Fig.
-10 (b): (1) without the last LN layer, the onset of discrimination shifts “forward".
-Specifically, the cosine similarity begins to change from the $8$th transformer block and exhibits a sharp decrease at the $10$th transformer block; (2) though the performance of our method (Pattern Retrieval) drops by $4.65\%$ without the final LN layer, it is still much higher that the baseline ($42.84\%$), demonstrating the robustness of our ImageStacker with pattern retrieval; (3) even when using the original image in conjunction with the query itself as the example pair (the upper bound), the $\mu AP$ does not remain at $100\%$.
-This indicates that the emergence of in-context learning is compromised to a certain degree.
-Figure 11: The dissection of the final LN layer.
-Figure 12: The influence of normalization and affine in the training stage.
-Diving into the final Layer Normalization layer. Given the input $x$, the LN is represented by:
-$y=\frac{x-\mathrm{E}[x]}{\sqrt{\operatorname{Var}[x]+\epsilon}}*\gamma+\beta,$ (6)
-where: $\epsilon$ is a value added to the denominator for numerical stability; $\gamma$ and $\beta$ are two learnable affine transform parameters.
-It can be dissected into normalization and affine.
-Fig.
-11 shows the change of similarity inside the LN layer.
-It is the affine transformation that separates different features.
-To delve deeper into the roles of normalization and affine transformation during training, we independently remove each: the normalization and the affine transformation.
-The modified versions of the final LN layer are represented as:
-$y_{1}=x*\gamma+\beta,$ (7)
-and
-$y_{2}=\frac{x-\mathrm{E}[x]}{\sqrt{\operatorname{Var}[x]+\epsilon}}.$ (8)
-The repeated experimental results are illustrated in Fig.
-12.
-We discover that: (1) as expected, removing either the normalization or the affine transformation results in a performance drop for our method ($-3.37\%$ and $-1.48\%$ respectively).
-However, both are still less than the drop observed when the last LN layer is entirely removed ($-4.65\%$); (2) the phenomenon in which discrimination shifts “forward" is attributed to the affine transformation.
-Specifically, without normalization, there is no “forward" shift, whereas without the affine transformation, a similar phenomenon is observed; (3) even without either the normalization or the affine transformation, the $\mu AP$ upper bound still remains near $100\%$.
-This indicates that this in-context learning phenomenon is not solely a result of the LN itself; even a basic affine transformation can produce a similar effect.
-The final Layer Normalization layer is also crucial for fundamental vision tasks like image classification.
-This section explores the influence of the final LN layer beyond our in-context ICD.
-We select the image classification task for demonstration due to its fundamental nature.
-The experiments are conducted with DeiT [46] on ImageNet [27], with results shown in Fig.
-13 and Table 4.
-We observe that the final LN layer plays a crucial role in (1) stabilizing the training process – without it, the training loss can fluctuate to as high as $35$; and in (2) enhancing the performance – for instance, in the absence of the final LN layer, the top-$1$ accuracy of DeiT-B drops by $2.4\%$ (from $81.8\%$ to $79.4\%$).
-These findings underscore the significance of the final LN layer, extending beyond the in-context ICD.
-Figure 13: The change of training loss of DeiT-B with respect to epochs.
-Table 4: Comparison of top-1 accuracy with and without the final LN layer.
-Top-1 Accuracy with last LN without last LN
-DeiT-Ti $72.2$ $68.4$
-DeiT-S $79.8$ $76.1$
-DeiT-B $81.8$ $79.4$
-Appendix D Evaluation metrics
-We employ two evaluation metrics for the in-context ICD, namely $\mu AP$ and $Recall@1$.
-$\mu AP$ serves as an overall evaluation metric and is equivalent to the area under the Precision-Recall curve, providing a comprehensive measure of both the precision and recall of our model across varying thresholds.
-On the other hand, the $Recall@1$ metric is query-specific.
-It checks whether the actual (correct) result appears first in the list of all returned results, offering insight into the effectiveness of our model in accurately retrieving the most relevant result at the top.
-Appendix E Training details
-We implement our ImageStacker using PyTorch [47] and distribute its training across eight Nvidia A100 GPUs.
-We use ViT-B/16 [24] as the backbone, which is pre-trained on the ImageNet dataset [27] using DeiT [46] unless otherwise specified.
-Before training, we resize images to a resolution of $224\times 224$ pixels.
-We set the balance parameter, $\lambda$, at $1$ and use a batch size of $512$.
-Each batch adopts the standard PK sampling method, with $128$ classes and $4$ images per class.
-The total number of training epochs is $25$ with a cosine-decreasing learning rate.
-The margin $m$ and scale $s$ in CosFace loss [22] are set to $0.35$ and $64$, respectively.
-Appendix F The performance on base patterns
-Table 5: The performance on the base patterns from AnyPattern and in-context learning.
-Our in-context learning method (ImageStacker) not only improves performance on novel patterns significantly but also maintains (marginally improves) the performance on base patterns.
-[…]
-AnyPattern 77.54 76.12 77.56 79.84 79.37
-ImageStacker 81.39 79.98 80.11 84.13 83.56
-[…]
-AnyPattern 81.76 80.37 81.13 83.41 83.05
-ImageStacker 84.52 83.96 83.78 86.07 85.85
-It is crucial to maintain performance on base patterns while enhancing performance on novel patterns through in-context learning because the edited copies generated by the base patterns may still appear in the future.
-As illustrated in Table 5, our proposed ImageStacker achieves this objective.
-For instance, with our baseline, the $\mu AP$ has been increased from $79.37\%$ to $83.56\%$, and the $recall@1$ has improved from $83.05\%$ to $85.85\%$.
-Appendix G The proposed pattern retrieval method is effective
-Table 6: Different methods to retrieve the image-replica pair of a query.
-[…]
-Basline (Any) $42.84$ $47.86$
-Random $50.42$ $54.83$
-Approximate $56.65$ $60.86$
-Accurate $56.68$ $60.78$
-Ground truth $56.99$ $60.96$
-Zero Shot $49.71$ $53.76$
-Lower Bound $28.55$ $32.25$
-Upper Bound $99.25$ $99.84$
-Figure 14: The demonstration of different methods for selecting an image-replica pair. A larger score means a better matching.
-Table 6 outlines several methods for obtaining the image-replica pair for a given query: (1) Random - selecting an image-replica pair randomly from the entire image-replica pool.
-(2) Approximate retrieval - using the ImageStacker model.
-(3) Accurate retrieval - employing another model.
-(4) Ground truth - using pattern ground truth of queries (not available in practice).
-ImageStacker can also be applied in a zero-shot setting by duplicating the query two times as the pseudo-prompt (Zero Shot).
-The lower bound is obtained by using an incorrect image-replica pair, while the upper bound is achieved by using the query and its original image as the image-replica.
-Their pattern retrieval accuracy is shown in the Appendix (Section H).
-The demonstration of these methods is visualized in the Fig.
-14 (See more demonstrations in Fig.
-17 and Fig.
-18).
-We observe that: (1) our approximate retrieval method achieves performance comparable to that of the accurate method ($-0.03\%$ $\mu AP$), and even comparable to using the pattern ground truth ($-0.34\%$ $\mu AP$).
-It significantly outperforms the random method ($+6.23\%$ $\mu AP$).
-(2) The zero-shot setting surpasses the baseline by $6.87\%$ in $\mu AP$ and $5.90\%$ in $recall@1$, demonstrating the generalizability of our method.
-(3) It is non-trivial that using the query itself and its original image as the image-replica nearly achieves $100\%$ performance (see Table 3: without our AnyPattern, using the query itself and its original image can only improve the $\mu AP$ from $16.18\%$ to $18.11\%$), providing evidence of the emergence of in-context learning again.
-Appendix H Our pattern retrieval method achieves high accuracy
-Table 7: The accuracy of pattern retrieval methods and corresponding $\mu AP$.
-[…]
-Random $30.11$ $50.42$
-Approximate $62.90$ $56.65$
-Accurate $64.60$ $56.68$
-Ground truth $100.00$ $56.99$
-This section shows the accuracy of our pattern retrieval methods achieve. The accuracy is defined as
-$acc\ =\ \frac{1}{N}\sum^{N}_{i=1}acc_{i},$ (9)
-$acc_{i}\ =\ \frac{\#\left(P_{q_{i}}\cap P_{s_{i}}\right)}{{}\#\left(P_{q_{i}}\right)},$ (10)
-[…]
-The pattern retrieval accuracy and the corresponding $\mu AP$ for the four retrieval methods are displayed in Table 7.
-Our observations are as follows: (1) compared to random selection, our methods of approximate and accurate pattern retrieval show an improvement of approximately $+30\%$ in accuracy, leading to a $+6\%$ improvement in $\mu AP$; (2) while using ground truths can achieve $100\%$ accuracy in pattern retrieval, there is little to no room for further improvement in $\mu AP$.
-These findings indicate that, although our pattern retrieval is not perfect, it is sufficient to retrieve an example pair to condition the feed-forward process.
-Appendix I Utilizing multiple image-replica pairs per query further improves the performance
-Figure 15: Leveraging multiple image-replica pairs per query: the change of $\mu AP$ and $recall@1$ in relation to the number of image-replica pairs per query.
-We also discover that leveraging multiple image-replica pairs for a given query can further enhance performance.
-For each query, we retrieve its top-$N$ image-replica pairs and repeat the $9$-channel ImageStacker $N$ times.
-The final similarity between a query and a gallery is then calculated as the maximum value among $N$ similarities.
-The performance change in relation to the number of image-replica pairs is depicted in Fig.
-15.
-Compared to using only one image-replica pair, employing $10$ image-replica pairs boosts the $\mu AP$ to $59.59\%$ ($+2.94\%$) and $recall@1$ to $63.16\%$ ($+2.30\%$).
-Appendix J Our method remains effective even with fewer image-replica pairs in the pool.
-Figure 16: The change of $\mu AP$ and $recall@1$ in relation to the number of image-replica pairs per pattern combination in the pool.
-Currently, the AnyPattern pool contains 10 image-replica pairs for each pattern combination.
-As shown in Fig.
-16, we find that for our ImageStacker, one image-replica pair per pattern combination is sufficient to improve performance.
-This implies that when we utilize only one image-replica pair per query for inference (unlike the scenario described in Section I), 10 image-replica pairs per pattern combination in the pool achieves similar performance to just one pair.
-This further demonstrates the practicality of our method: when operators of a deployed ICD system identify novel patterns, they need only manually add one example pair to the pool.
-Appendix K Author statement
-We, the undersigned, hereby affirm that we bear full responsibility for the content provided herein.
-We guarantee that all data and materials used and presented are compliant with applicable laws and are free from any encroachments on intellectual property rights or any other legal rights of third parties.
-In the event of any violations, we accept full liability and will address all associated legal consequences.
-Appendix L Hosting, licensing, and maintenance plan.
-The data is hosted by Hugging Face, Inc.
-at https://huggingface.co/datasets/WenhaoWang/AnyPattern.
-The dataset and corresponding codes are released under the MIT License.
-We commit to maintaining the dataset, which includes regular updates and corrections as necessary.
-Figure 17: More visualizations of different methods for selecting an example pair. A larger score means a better matching.
-Figure 18: More visualizations of different methods for selecting an example pair. A larger score means a better matching.
-Appendix M Data Sheet for AnyPattern
+Appendix B Data Sheet for AnyPattern
 For what purpose was the dataset created? Was there a specific task in mind? Was there a specific gap that needed to be filled? Please provide a description.
@@
 Who created the dataset (e.g., which team, research group) and on behalf of which entity (e.g., company, institution, organization)?
-The dataset was created by Wenhao Wang (University of Technology Sydney), Yifan Sun (Baidu Inc.), Zhentao Tan (Peking University and Baidu Inc.), and Yi Yang (Zhejiang University).
+The dataset was created by Wenhao Wang (University of Technology Sydney), Yifan Sun (Baidu Inc.), Zhentao Tan (Baidu Inc.), and Yi Yang (Zhejiang University).
 Who funded the creation of the dataset? If there is an associated grant, please provide the name of the grantor and the grant name and number.
@@
 None.
-NeurIPS Paper Checklist
-  1. 1.
-    Claims
-  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
-  3. Answer: [Yes]
-  4. Justification: We follow this common practice for writing a paper.
-  5. Guidelines:
-    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
-    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
-A No or NA answer to this question will not be perceived well by the reviewers.
-    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
-    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
-  6. 2.
-    Limitations
-  7. Question: Does the paper discuss the limitations of the work performed by the authors?
-  8. Answer: [Yes]
-  9. Justification: See the Limitation at the end of our main paper.
-  10. Guidelines:
-    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
-    • The authors are encouraged to create a separate "Limitations" section in their paper.
-    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
-The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
-    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
-In general, empirical results often depend on implicit assumptions, which should be articulated.
-    • The authors should reflect on the factors that influence the performance of the approach.
-For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
-Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
-    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
-    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
-    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
-The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
-Reviewers will be specifically instructed to not penalize honesty concerning limitations.
-  11. 3.
-    Theory Assumptions and Proofs
-  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
-  13. Answer: [N/A]
-  14. Justification: This is not a theory paper.
-  15. Guidelines:
-    • The answer NA means that the paper does not include theoretical results.
-    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
-    • All assumptions should be clearly stated or referenced in the statement of any theorems.
-    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
-    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
-    • Theorems and Lemmas that the proof relies upon should be properly referenced.
-  16. 4.
-    Experimental Result Reproducibility
-  17.
-Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
-  18. Answer: [Yes]
-  19. Justification: The project (including the proposed dataset AnyPattern and the code for ImageStacker) is publicly available at https://anypattern.github.io under the MIT Licence.
-  20. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
-    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
-    • Depending on the contribution, reproducibility can be accomplished in various ways.
-For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
-In general.
-releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
-    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
-For example
-      1. (a)
-        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
-      2. (b)
-        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
-      3. (c)
-        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
-      4. (d)
-        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
-In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
-  21. 5.
-    Open access to data and code
-  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
-  23. Answer: [Yes]
-  24. Justification: The project (including the proposed dataset AnyPattern and the code for ImageStacker) is publicly available at https://anypattern.github.io under the MIT Licence.
-  25. Guidelines:
-    • The answer NA means that paper does not include experiments requiring code.
-    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
-    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
-Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
-    • The instructions should contain the exact command and environment needed to run to reproduce the results.
-See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
-    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
-    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
-If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
-    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
-    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
-  26. 6.
-    Experimental Setting/Details
-  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
-  28. Answer: [Yes]
-  29. Justification: The experimental settings and details are provided in the main paper as well as the Appendix.
-  30. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
-    • The full details can be provided either with the code, in appendix, or as supplemental material.
-  31. 7.
-    Experiment Statistical Significance
-  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
-  33. Answer: [No]
-  34. Justification: All the experimental results are significant and stable, and error bars are not reported because it would be too computationally expensive.
-  35. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
-    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
-    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
-    • The assumptions made should be given (e.g., Normally distributed errors).
-    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
-    • It is OK to report 1-sigma error bars, but one should state it.
-The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
-    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
-    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
-  36. 8.
-    Experiments Compute Resources
-  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
-  38. Answer: [Yes]
-  39. Justification: We include the experiments compute resources in each related section.
-  40. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
-    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
-    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
-  41. 9.
-    Code Of Ethics
-  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
-  43. Answer: [Yes]
-  44. Justification: We follow the NeurIPS Code of Ethics.
-  45. Guidelines:
-    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
-    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
-    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
-  46. 10.
-    Broader Impacts
-  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
-  48. Answer: [Yes]
-  49. Justification: Please see the end of Conclusion and the Data Sheet.
-  50. Guidelines:
-    • The answer NA means that there is no societal impact of the work performed.
-    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
-    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
-    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
-However, if there is a direct path to any negative applications, the authors should point it out.
-For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
-On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
-    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
-    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
-  51. 11.
-    Safeguards
-  52.
-Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
-  53. Answer: [N/A]
-  54. Justification: Our paper poses no such risks.
-  55. Guidelines:
-    • The answer NA means that the paper poses no such risks.
-    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
-    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
-    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
-  56. 12.
-    Licenses for existing assets
-  57.
-Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
-  58. Answer: [Yes]
-  59. Justification: We cite all the assets used in our paper, and all the corresponding licenses are mentioned and respected.
-  60. Guidelines:
-    • The answer NA means that the paper does not use existing assets.
-    • The authors should cite the original paper that produced the code package or dataset.
-    • The authors should state which version of the asset is used and, if possible, include a URL.
-    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
-    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
-    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
-For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
-Their licensing guide can help determine the license of a dataset.
-    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
-    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
-  61. 13.
-    New Assets
-  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
-  63. Answer: [Yes]
-  64. Justification: The project (including the proposed dataset AnyPattern and the code for ImageStacker) is publicly available at https://anypattern.github.io under the MIT Licence.
-  65. Guidelines:
-    • The answer NA means that the paper does not release new assets.
-    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
-    • The paper should discuss whether and how consent was obtained from people whose asset is used.
-    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
-  66. 14.
-    Crowdsourcing and Research with Human Subjects
-  67.
-Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
-  68. Answer: [N/A]
-  69. Justification: The paper does not involve crowdsourcing nor research with human subjects.
-  70. Guidelines:
-    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
-    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
-    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
-  71. 15.
-    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
-  72.
-Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
-  73. Answer: [N/A]
-  74. Justification: We do not involve research with human subjects and there is no potential risk.
-  75. Guidelines:
-    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
-    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
-If you obtained IRB approval, you should clearly state this in the paper.
-    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
-    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.